### v1

# Nash Equilibrium between Brokers and Traders

Abstract
We study the perfect information Nash equilibrium between a broker and her clients — an informed trader, and an uniformed trader. In our model, the broker trades in the lit exchange where trades have instantaneous and transient price impact with exponential resilience, while both clients trade with the broker. The informed trader and the broker maximise expected wealth subject to inventory penalties, while the uninformed trader is not strategic and sends the broker random buy and sell orders. We characterise the Nash equilibrium of the trading strategies with the solution to a coupled system of forward-backward stochastic differential equations (FBSDEs). We solve this system explicitly and study the effect of information in the trading strategies of the broker and the informed trader.
journal: TBA
1 Introduction
This paper characterises the Nash equilibrium between a broker and her clients in an over-the-counter (OTC) market. The problem formulation is as follows. The broker provides liquidity to both an uninformed trader and an informed trader, and the broker also trades in a lit exchange where her buy and sell market orders have transient and instantaneous price impact. The trading flow of the uniformed trader is an exogenous process that mean-reverts around zero. The informed trader knows the stochastic drift of the asset and trades strategically. On the other hand, the broker knows the identity of her clients and uses this information to devise an internalisation and externalisation strategy of her inventory. In particular, the broker’s strategy uses the flow of the informed trader to extract information about the trend in the price of the asset, which she uses to hedge her inventory position and to execute speculative trades in the lit exchange.
The setup of the interaction between the traders and the broker is similar to that in Cartea and Sánchez-Betancourt, (2022). Specifically, the broker and the informed trader maximise expected wealth from their trading activities, while penalising inventory holdings. For the broker, this penalty protects her strategy from inventory risk, in particular toxic inventory. Similarly, for the informed trader, the inventory penalty controls how much inventory risk he is willing to bear throughout the trading horizon. In Cartea and Sánchez-Betancourt, (2022), the broker’s trades in the lit market have a permanent impact on prices, whereas here the broker’s trades have both instantaneous and transient impact (see e.g., Neuman and Voß, (2022)), where the latter is modelled with exponential kernels as in Obizhaeva and Wang, (2013).
In our model, the broker controls her speed of trading $\nu$ in the lit exchange and the informed trader controls the speed $\eta$ at which he trades with the broker. The performance criteria of the broker and of the informed trader are $J^{B}(\nu,\eta)$ and $J^{I}(\nu,\eta)$, respectively. We show that for a large set of admissible controls: (i) for fixed trading speed $\nu$ of the broker, the map $\eta\to J^{I}(\nu,\eta)$ is strictly concave, and (ii) for fixed trading speed $\eta$ of the informed trader, the map $\nu\to J^{B}(\nu,\eta)$ is strictly concave. We use this result and the Gâteaux derivative of the functionals $J^{I},J^{B}$ to characterise the best responses of the broker and the informed trader. Then, we show that the Nash equilibrium of the strategies of the broker and of the informed trader is given by the solution to a linear forward-backward stochastic differential equation (FBSDE). We show that there is a unique solution to the FBSDE that characterises the Nash equilibrium if and only if there is a solution to a matrix Riccati differential equation. We show existence and uniqueness of the matrix Riccati differential equation under some restrictions of model parameters. We also present a proof of existence and uniqueness at the level of the FBSDE with an explicit bound on the time horizon. Finally, we use the closed-form Nash equilibrium strategies to perform simulations and showcase the performance of the trading strategies. We compare the performance of the strategies in the Nash equilibrium with the strategies in the two-stage optimisation of Cartea and Sánchez-Betancourt, (2022) and derive insights.
There are a number of studies that focus on the equilibrium between takers and makers of liquidity in financial markets. Earlier works include those of Grossman and Stiglitz, (1980), Kyle, (1985), Kyle, (1989), and O’Hara, (1998). Grossman and Miller, (1988) study the liquidity of a traded asset as a result of the demand and supply of immediacy in a market where liquidity providers charge a premium for intermediating trades between two consecutive trading periods. Recently, Bank et al., (2021) extend the Grossman–Miller model to a continuous-time framework where a risky asset is traded in two markets, a lit exchange, and an OTC dealer market. Similar to Bank et al., our work considers a broker who decides how to internalise and externalise trades. While in their model, “pure internalisers only trade in the dealer market, whereas externalisation corresponds to offsetting trades in the open market”, see also Butz and Oomen, (2019), in our model, the broker’s strategy consists of internalising and externalising trades, and of speculative trades that are based on the information learned from the flow of the informed trader.
Herdegen et al., (2023) study a game between dealers that compete for the order flow of a client. Other stochastic games in the algorithmic trading literature study the interactions between agents liquidating inventory positions; see e.g., Cont et al., (2023); Abi Jaber et al., (2023). The interaction between the broker and traders in this paper departs from this branch of the literature. Here, the incentives to trade are as follows: the informed trader profits from the private signal about the trend in the value of the asset, the uninformed trader is not strategic, and the broker manages inventory and uses the trading speed of the informed trader to send speculative trades to the lit market. [^1] This is in contrast to the work of Cartea et al., (2023) where machine learning techniques determine if individual trades are informed or uninformed.
Another branch in the literature focuses on how a trader unwinds stochastic order flow. Cartea et al., (2020) derive an optimal liquidation strategy for a broker who makes liquidity and trades in a triplet of currency pairs, while managing stochastic order flow from her clients, in the presence of model uncertainty. In Cartea et al., (2022) the stochastic order flow is the proceeds from trading a stock in a foreign currency, so the broker must also manage exchange rate risk. Nutz et al., (2023) solve a control problem for the optimal externalisation schedule of an exogenous order flow with an Obizhaeva–Wang-type price impact and quadratic instantaneous costs. Recently, Barzykin et al., (2024) employ stochastic filtering to solve the control problem of unwinding stochastic order flow with unobserved toxicity. In our paper, the broker’s objective is to manage the stochastic order flow she receives from her clients; the optimal strategies are those that balance hedging, speculative trading, and inventory control; similar to the previous two works, we also employ Obizhaeva–Wang-type price impact with quadratic instantaneous costs.
Finally, in a mean field setup, Baldacci et al., (2023) study a model with one market maker and the average behaviour of an infinite number of liquidity takers. In a similar vein, Bergault and Sánchez-Betancourt, (2024) characterise the mean-field Nash equilibrium of the strategies between a broker, whose trades have instantaneous and permanent price impact, and a large number of informed traders. In this paper we study the optimal responses of a broker and and informed trader, which represents a collection of informed traders, where the broker’s trades have instantaneous and transient price impact with exponential resilience.
The remainder of the paper proceeds as follows. Section 2 introduces the model, the performance criteria, and defines the notion of a Nash equilibrium in our setup. Section 3 characterises the Nash equilibrium of the game and solves it explicitly. First, we show that both functionals are strictly concave (up to negligible sets), then, we compute their Gâteaux derivatives, we characterise the Nash equilibrium with a system of FBSDEs, and we study existence and uniqueness of the system. Section 4 derives the closed-form solution and connects the existence and uniqueness of the FBSDE with the existence and uniqueness of a matrix Riccati differential equation; we show existence and uniqueness of this differential equation under some restrictions of model parameters. Finally, Section 5 shows numerical experiments and concludes.
2 The model
Let $T>0$ be a given trading horizon and let ${\mathfrak{T}}=[0,T]$. We fix a filtered probability space $\left(\Omega,\allowbreak{\mathcal{F}},({\mathcal{F}}_{t})_{t\in{\mathfrak{T}}},\mathbb{P}\right)$ satisfying the usual conditions of right-continuity and completeness. The filtration $({\mathcal{F}}_{t})_{t\in{\mathfrak{T}}}$ is the sigma-algebra generated by all processes below.
The broker trades with speed $(\nu_{t})_{t\in{\mathfrak{T}}}$, the informed trader trades with speed $(\eta_{t})_{t\in{\mathfrak{T}}}$, and the uninformed trader trades with speed $(\xi_{t})_{t\in{\mathfrak{T}}}$. We assume that $\nu,\eta,\xi\in{\mathbb{H}^{2}}$ where
${\mathbb{H}^{2}}=\Big{\{}(\gamma_{t})_{t\in{\mathfrak{T}}}\,\,|\,\,\gamma\text{ is }{\mathcal{F}}-\text{progressively measurable, and }\|\gamma\|^{2}:=\mathbb{E}\left[\textstyle\int_{0}^{T}\left(\eta_{s}\right)^{2}\mathrm{d}s\right]\,<\,+\infty\Big{\}}\,.$ (2.1)
The midprice process is denoted by $(S_{t})_{t\in{\mathfrak{T}}}$ and is given by
$S_{t}=S_{0}+\underbrace{\int_{0}^{t}\alpha_{s}\,\mathrm{d}s}_{\text{signal}}+\underbrace{Y_{t}}_{\text{impact}}+M^{S}_{t}\,,$ (2.2)
where $(M^{S}_{t})_{t\in{\mathfrak{T}}}$ is a square-integrable martingale with respect to $({\mathcal{F}}_{t})_{t\in{\mathfrak{T}}}$, the broker’s price price impact $(Y_{t})_{t\in{\mathfrak{T}}}$ satisfies
$\mathrm{d}Y_{t}=(\mathfrak{h}\,\nu_{t}-\mathfrak{p}\,Y_{t})\,\mathrm{d}t\,,\quad Y_{0}\in\mathbb{R}\,,$ (2.3)
and the trading signal process $(\alpha_{t})_{t\in{\mathfrak{T}}}$ is square-integrable and progressively measurable. Here, $\mathfrak{p}\geq 0$ is a decay coefficient, and $\mathfrak{h}\geq 0$ is the instantaneous price impact parameter. When $\mathfrak{p}=0$ the price impact is permanent and linear in the integral of the broker’s speed of trading.
The inventory process $Q^{B}$ and cash process $X^{B}$ of the broker satisfy the equations
$\mathrm{d}Q^{B}_{t}$ $=\left(\nu_{t}-\eta_{t}-\xi_{t}\right)\mathrm{d}t\,,$ $Q^{B}_{0}=q^{B}\in\mathbb{R}\,,$ (2.4)
$\mathrm{d}X^{B}_{t}$ $=-\left(S_{t}+a\,\nu_{t}\right)\nu_{t}\,\mathrm{d}t+\left(S_{t}+b\,\eta_{t}\right)\eta_{t}\,\mathrm{d}t+\left(S_{t}+c\,\xi_{t}\right)\xi_{t}\,\mathrm{d}t\,,$ $X^{B}_{0}=0\,,$ (2.5)
where the positive constants $a,b,c$ represent instantaneous transaction costs. We assume that the trading speed of the uninformed trader is square-integrable. Similarly, the inventory process $Q^{I}$ and cash process $X^{I}$ of the informed trader satisfy the equations
$\mathrm{d}Q^{I}_{t}$ $=\eta_{t}\,\mathrm{d}t\,,$ $Q^{I}_{0}=q^{I}\in\mathbb{R}\,,$ (2.6)
$\mathrm{d}X^{I}_{t}$ $=-\left(S_{t}+b\,\eta_{t}\right)\eta_{t}\,\mathrm{d}t\,,$ $X^{I}_{0}=0\,.$ (2.7)
Sometimes we write the control in the superscript when we highlight the controls that affect a process, for example, we write $X^{I,\eta}$ and $Q^{I,\eta}$ for the cash process and inventory process of the informed trader. The sets of admissible strategies for the informed trader and the broker are both given by the set ${\mathbb{H}^{2}}$ of squared-integrable adapted processes.
Let $\nu\in{\mathbb{H}^{2}}$ and let $\eta\in{\mathbb{H}^{2}}$. The performance criterion of the informed trader is $J^{I}(\nu,\eta)$, given by
$J^{I}(\nu,\eta)=\mathbb{E}\left[X^{I}_{T}+Q^{I}_{T}\,S_{T}-\psi\,\left(Q^{I}_{T}\right)^{2}-r^{I}\int_{0}^{T}\left(Q^{I}_{s}\right)^{2}\,\mathrm{d}s\right]\,,$ (2.8)
for $\psi,r^{I}$ non-negative inventory control constants. Similarly, the performance criterion of the broker is $J^{B}(\nu,\eta)$, given by
$J^{B}(\nu,\eta)=\mathbb{E}\left[X^{B}_{T}+Q^{B}_{T}\,S_{T}-\phi\,\left(Q^{B}_{T}\right)^{2}-r^{B}\int_{0}^{T}\left(Q^{B}_{s}\right)^{2}\,\mathrm{d}s\right]\,,$ (2.9)
for $\phi,r^{B}$ non-negative inventory control constants. We let
$\varphi=\phi-\frac{1}{2}\,\mathfrak{h}\,,$ (2.10)
and we assume that $\varphi\geq 0$ and that $a>\mathfrak{p}\,\mathfrak{h}\,T$. These are non-restrictive technical assumptions which are sufficient conditions to guarantee strict concavity of $J^{B}$ up to null sets, and that we also use to prove existence of a non-symmetric matrix Riccati differential equation that arises in the closed-form solution to Nash equilibrium of the problem. In particular, the assumption that $a>\mathfrak{p}\,\mathfrak{h}\,T$ is not restrictive because one expects $a\gg\mathfrak{h}\,\mathfrak{p}$; see, e.g., page 148 in Cartea et al., (2015).
Lemma 2.1 (Finiteness of Performance Criterion).
The functional $J^{I}:{\mathbb{H}^{2}}\times{\mathbb{H}^{2}}\to\mathbb{R}$ can be written as
$J^{I}(\nu,\eta)=S_{0}\,q^{I}-\psi\left(q^{I}\right)^{2}+\mathbb{E}\left[\int_{0}^{T}\left\{-b\,\eta^{2}_{t}+Q^{I}_{t}\,\Big{(}\alpha_{t}+(\mathfrak{h}\,\nu_{t}-\mathfrak{p}\,Y_{t})-2\,\psi\,\eta_{t}-r^{I}\,Q^{I}_{t}\Big{)}\right\}\mathrm{d}t\right].$ (2.11)
Similarly, the functional $J^{B}:{\mathbb{H}^{2}}\times{\mathbb{H}^{2}}\to\mathbb{R}$ can be written as
$J^{B}(\nu,\eta)$ $=S_{0}\,q^{B}-\phi\,\left(q^{B}\right)^{2}$ (2.12)
$\quad+\mathbb{E}\left[\int_{0}^{T}\left\{-a\,\nu^{2}_{t}+b\,\eta^{2}_{t}+c\,\xi^{2}_{t}+Q^{B}_{t}\,\Big{(}\alpha_{t}+(\mathfrak{h}\,\nu_{t}-\mathfrak{p}\,Y_{t})-2\,\phi\,(\nu_{t}-\eta_{t}-\xi_{t})-r^{B}\,Q^{B}_{t}\Big{)}\right\}\mathrm{d}t\right].$
Proof.
The representation follows immediately from the product rule and the dynamics above. To see that for $\eta\in{\mathbb{H}^{2}}$ and $\nu\in{\mathbb{H}^{2}}$ both $J^{I}(\nu,\eta)\in\mathbb{R}$ and $J^{B}(\nu,\eta)\in\mathbb{R}$ (i.e., the performance criterion are finite), we proceed as follows. Use the inequality $2\,|x\,y|\leq x^{2}+y^{2}$, together with the following three results:
  1. (i)
    if $\eta\in{\mathbb{H}^{2}}$, then
    $\left(Q^{I,\eta}_{t}\right)^{2}$ $=\left(q_{0}^{I}\right)^{2}+2\,q_{0}^{I}\,\int_{0}^{t}\eta_{s}\,\mathrm{d}s+\left(\int_{0}^{t}\eta_{s}\,\mathrm{d}s\right)^{2}$ (2.13)
    $\leq\left(q_{0}^{I}\right)^{2}+\left|q_{0}^{I}\right|\,\int_{0}^{t}\left(1+\eta^{2}_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\left(\eta_{s}\right)^{2}\mathrm{d}s\,,$ (2.14)
    thus, there exists $C\in\mathbb{R}^{+}$ such that
    $\left|Q^{I,\eta}_{t}\right|^{2}\leq C\left(1+\int_{0}^{t}|\eta_{s}|^{2}\,\mathrm{d}s\right)\,,$ (2.15)
    which implies that
    $\int_{0}^{T}\left|Q^{I,\eta}_{t}\right|^{2}\,\mathrm{d}t\leq C\,T\left(1+\int_{0}^{T}|\eta_{s}|^{2}\,\mathrm{d}s\right)<+\infty\,.$ (2.16)
    because $\eta\in{\mathbb{H}^{2}}$.
  2. (ii)
    if $\eta\in{\mathbb{H}^{2}}$, $\nu\in{\mathbb{H}^{2}}$, and $\xi$ is square-integrable, similar to the above, there exists $C\in\mathbb{R}^{+}$ such that
    $\int_{0}^{T}\left|Q^{B,\nu}_{t}\right|^{2}\,\mathrm{d}t\leq C\,T\left(1+\left(\int_{0}^{T}|\eta_{s}|^{2}\,\mathrm{d}s+\int_{0}^{T}|\nu_{s}|^{2}\,\mathrm{d}s+\int_{0}^{T}|\xi_{s}|^{2}\,\mathrm{d}s\right)\right)<+\infty\,.$ (2.17)
  3. (iii)
    if $\nu\in{\mathbb{H}^{2}}$, then, from (2.3),
    $Y_{t}=e^{-\mathfrak{p}\,t}\,Y_{0}+\mathfrak{h}\,\int_{0}^{t}\nu_{s}\,e^{-\mathfrak{p}\,(t-s)}\,\mathrm{d}s\,.$ (2.18)
    Then, there exists $C\in\mathbb{R}^{+}$ such that
    $\int_{0}^{T}|Y_{t}|^{2}\mathrm{d}t\leq C\,T\left(1+\int_{0}^{t}|\nu_{s}|^{2}\,\mathrm{d}s\right)<+\infty.$ (2.19)
Finally, using that $\eta,\xi,\nu,Q^{I,\eta},Q^{B,\nu},\alpha,Y$ are all square-integrable, together with the representations in (2.11) and (2.12) concludes the proof. ∎
Definition 2.2.
The pair of trading speeds $\left(\nu^{*},\eta^{*}\right)\in{\mathbb{H}^{2}}\times{\mathbb{H}^{2}}$ is a Nash equilibrium of the strategies between the informed trader and the broker if the following two conditions hold:
  1. (i)
    for any speed of trading $\nu\in{\mathbb{H}^{2}}$
    $J^{B}(\nu,\eta^{*})\leq J^{B}(\nu^{*},\eta^{*})\,,$ (2.20)
  2. (ii)
    for any speed of trading $\eta\in{\mathbb{H}^{2}}$
    $J^{I}(\nu^{*},\eta)\leq J^{I}(\nu^{*},\eta^{*})\,.$ (2.21)
In the next section we characterise the perfect information Nash equilibrium of the trading strategies.
3 Nash equilibrium
3.1 Characterisation of equilibrium
We employ Gâteaux derivatives to characterise the Nash equilibrium between the broker and the informed trader. Let $\nu,\nu^{\prime}\in{\mathbb{H}^{2}}$ and $\eta,\eta^{\prime}\in{\mathbb{H}^{2}}$. The directional derivative of $J^{B}$ at $\nu$ in the direction of $\mathfrak{v}:=(\nu^{\prime}-\nu)$ is given by
$\langle\mathcal{D}\,J^{B}(\nu,\eta),\mathfrak{v}\rangle=\lim_{\varepsilon\to 0}\tfrac{1}{\varepsilon}\left[J^{B}(\nu+\varepsilon\,\mathfrak{v},\eta)-J^{B}(\nu,\eta)\right]\,,$ (3.1)
when the limit exists. Similarly, the directional derivative of $J^{I}$ at $\eta$ in the direction of $\mathfrak{n}:=(\eta^{\prime}-\eta)$ is given by
$\langle\mathcal{D}\,J^{I}(\nu,\eta),\mathfrak{n}\rangle=\lim_{\varepsilon\to 0}\tfrac{1}{\varepsilon}\left[J^{I}(\nu,\eta+\varepsilon\,\mathfrak{n})-J^{I}(\nu,\eta)\right]\,,$ (3.2)
when the limit exists.
First, we show that for fixed $\nu\in{\mathbb{H}^{2}}$ the functional $\eta\to J^{I}(\nu,\eta)$ is strictly concave up to negligible sets; this is proven in the next proposition.
Proposition 3.1.
Let $\nu\in{\mathbb{H}^{2}}$, the functional $J^{I}(\nu,\,\,\cdot\,):{\mathbb{H}^{2}}\to\mathbb{R}$ is strictly concave.
Proof.
Let $\nu\in{\mathbb{H}^{2}}$ and $\eta,\kappa\in{\mathbb{H}^{2}}$. Let $A\in{\mathcal{F}}\otimes{\mathcal{B}}({\mathfrak{T}})$ with $\mu(A)>0$, where $\mu:=\mathbb{P}\otimes\mathrm{d}t$ and for all $(\omega,t)\in A$ we have that $\eta_{t}(\omega)\neq\kappa_{t}(\omega)$, i.e., $\eta$ and $\kappa$ differ on a set with non-zero $\mu$-measure. Let $\rho\in(0,1)$, we need to show that
$J^{I}(\nu,\rho\,\eta+(1-\rho)\,\kappa)>\rho\,J^{I}(\nu,\eta)+(1-\rho)\,J^{I}(\nu,\kappa)\,.$ (3.3)
First, from linearity of (2.6), observe that
$Q^{I,\rho\,\eta+(1-\rho)\,\kappa}_{t}=\rho\,Q^{I,\eta}_{t}+(1-\rho)\,Q^{I,\kappa}_{t}\,,$ (3.4)
thus,
$J^{I}(\nu,\rho\,\eta+(1-\rho)\,\kappa)$ (3.5)
$\quad=S_{0}\,q^{I}-\psi\left(q^{I}\right)^{2}+\mathbb{E}\Bigg{[}\int_{0}^{T}\bigg{\{}-b\,\left(\rho\,\eta_{t}+(1-\rho)\,\kappa_{t}\right)^{2}+\left(\rho\,Q^{I,\eta}_{t}+(1-\rho)\,Q^{I,\kappa}_{t}\right)\,(\alpha_{t}+(\mathfrak{h}\,\nu_{t}-\mathfrak{p}\,Y_{t}))$
$\hskip 140.00021pt-2\,\psi\,\left(\rho\,Q^{I,\eta}_{t}+(1-\rho)\,Q^{I,\kappa}_{t}\right)\,\left(\rho\,\eta_{t}+(1-\rho)\,\kappa_{t}\right)$
$\hskip 140.00021pt-r^{I}\,\left(\rho\,Q^{I,\eta}_{t}+(1-\rho)\,Q^{I,\kappa}_{t}\right)^{2}\bigg{\}}\,\mathrm{d}t\Bigg{]}$ (3.6)
$\quad=\rho\,J^{I}(\nu,\eta)+\rho\,(1-\rho)\,\mathbb{E}\left[\int_{0}^{T}\left\{b\,\eta^{2}_{t}+2\,\psi\,Q^{I,\eta}\,\eta_{t}+r^{I}\,\left(Q^{I,\eta}_{t}\right)^{2}\right\}\mathrm{d}t\right]$
$\hskip 62.50009pt+(1-\rho)\,J^{I}(\nu,\kappa)+\rho\,(1-\rho)\mathbb{E}\left[\int_{0}^{T}\left\{b\,\kappa^{2}_{t}+2\,\psi\,Q^{I,\kappa}\,\kappa_{t}+r^{I}\,\left(Q^{I,\kappa}_{t}\right)^{2}\right\}\mathrm{d}t\right]$
$\hskip 62.50009pt-2\,\rho\,(1-\rho)\,\mathbb{E}\left[\int_{0}^{T}\left\{b\,\eta_{t}\,\kappa_{t}+\psi\,Q^{I,\eta}\,\kappa_{t}+\psi\,Q^{I,\kappa}\,\eta_{t}+r^{I}\,Q^{I,\kappa}_{t}\,Q^{I,\eta}_{t}\right\}\mathrm{d}t\right]$ (3.7)
$\quad=\rho\,J^{I}(\nu,\eta)+(1-\rho)\,J^{I}(\nu,\kappa)$
$\qquad\qquad+\rho\,(1-\rho)\,\mathbb{E}\left[\int_{0}^{T}\left\{b\left(\eta_{t}-\kappa_{t}\right)^{2}+2\,\psi\left(Q^{\eta}_{t}-Q^{\kappa}_{t}\right)\left(\eta_{t}-\kappa_{t}\right)+r^{I}\,\left(Q^{I,\eta}_{t}-Q^{I,\kappa}_{t}\right)^{2}\right\}\mathrm{d}t\right]\,.$ (3.8)
Then, by the definition of $Q^{I}$ we have
$\mathbb{E}\left[\int_{0}^{T}\left(Q^{I,\eta}_{t}-Q^{I,\kappa}_{t}\right)\left(\eta_{t}-\kappa_{t}\right)\mathrm{d}t\right]$ $=\mathbb{E}\left[\int_{0}^{T}\left(\int_{0}^{t}(\eta_{u}-\kappa_{u})\mathrm{d}u\right)\left(\eta_{t}-\kappa_{t}\right)\mathrm{d}t\right]$
$=\frac{1}{2}\,\mathbb{E}\left[\int_{0}^{T}\int_{0}^{T}\left(\eta_{u}-\kappa_{u}\right)\left(\eta_{t}-\kappa_{t}\right)\mathrm{d}u\,\mathrm{d}t\right]$
$=\frac{1}{2}\,\mathbb{E}\left[\left(\int_{0}^{T}(\eta_{t}-\kappa_{t})\,\mathrm{d}t\right)^{2}\right]\,,$ (3.9)
where the equality in the second line follows because $\eta,\kappa\in{\mathbb{H}^{2}}$, hence, we apply Fubini’s theorem. Therefore,
$J^{I}(\nu,\rho\,\eta+(1-\rho)\,\kappa)=$ $\;\;\rho\,J^{I}(\nu,\eta)+(1-\rho)\,J^{I}(\nu,\kappa)$
$+\rho\,(1-\rho)\,\mathbb{E}\left[\int_{0}^{T}\left\{b\left(\eta_{t}-\kappa_{t}\right)^{2}+r^{I}\,\left(Q^{I,\eta}_{t}-Q^{I,\kappa}_{t}\right)^{2}\right\}\mathrm{d}t\right]$
$+\rho\,(1-\rho)\,\psi\,\mathbb{E}\left[\left(\int_{0}^{T}\eta_{t}-\kappa_{t}\,\mathrm{d}t\right)^{2}\right]\,.$
Hence, $J^{I}(\nu,\rho\,\eta+(1-\rho)\,\kappa)>\rho\,J^{I}(\nu,\eta)+(1-\rho)\,J^{I}(\nu,\kappa)$ because $\mu(A)>0$ and $\mathbb{E}\left[\int_{0}^{T}\left(\eta_{t}-\kappa_{t}\right)^{2}\mathrm{d}t\right]>0$. ∎
The proposition below shows that for fixed $\eta\in{\mathbb{H}^{2}}$ the broker’s functional $\nu\to J^{B}(\eta,\nu)$ is also strictly concave.
Proposition 3.2.
Let $\eta\in{\mathbb{H}^{2}}$, the functional $J^{B}(\,\cdot\,,\,\eta):{\mathbb{H}^{2}}\to\mathbb{R}$ is strictly concave.
Proof.
Let $\eta\in{\mathbb{H}^{2}}$ and $\nu,\zeta\in{\mathbb{H}^{2}}$. Let $A\in{\mathcal{F}}\otimes{\mathcal{B}}({\mathfrak{T}})$ with $\mu(A)>0$, where $\mu:=\mathbb{P}\otimes\mathrm{d}t$, and for all $(\omega,t)\in A$ we have that $\nu_{t}(\omega)\neq\zeta_{t}(\omega)$, i.e., $\nu$ and $\zeta$ differ on a set with non-zero $\mu$-measure. Let $\rho\in(0,1)$, we need to show that
$J^{B}(\rho\,\nu+(1-\rho)\,\zeta,\eta)>\rho\,J^{B}(\nu,\eta)+(1-\rho)\,J^{B}(\zeta,\eta)\,.$ (3.10)
First, from linearity of (2.4), observe that
$Q^{B,\rho\,\nu+(1-\rho)\,\zeta}_{t}=\rho\,Q^{B,\nu}_{t}+(1-\rho)\,Q^{B,\zeta}_{t}$ (3.11)
and, from linearity of (2.3),
$Y^{\rho\,\nu+(1-\rho)\,\zeta}_{t}=e^{-\mathfrak{p}\,t}\,Y_{0}+\mathfrak{h}\,\int_{0}^{t}e^{-\mathfrak{p}\,(t-s)}\,\left(\rho\,\nu_{s}+(1-\rho)\,\zeta_{s}\right)\,\mathrm{d}s=\rho\,Y^{\nu}_{t}+(1-\rho)\,Y^{\zeta}_{t}\,.$ (3.12)
Next,
$J^{B}(\rho\,\nu+(1-\rho)\,\zeta,\eta)=S_{0}\,q^{B}-\phi\,\left(q^{B}\right)^{2}$ (3.13)
$\;+\mathbb{E}\Bigg{[}\int_{0}^{T}\bigg{\{}-a\,\left(\rho\,\nu_{t}+(1-\rho)\,\zeta_{t}\right)^{2}+b\,\eta^{2}_{t}+c\,\xi^{2}_{t}+\left(\rho\,Q^{B,\nu}_{t}+(1-\rho)\,Q^{B,\zeta}_{t}\right)\,\alpha_{t}$
$\quad\qquad+\left(\rho\,Q^{B,\nu}_{t}+(1-\rho)\,Q^{B,\zeta}_{t}\right)\,\left(\mathfrak{h}\,\rho\,\nu_{t}+\mathfrak{h}\,(1-\rho)\,\zeta_{t}-\mathfrak{p}\,\rho\,Y^{\nu}_{t}-\mathfrak{p}\,(1-\rho)\,Y^{\zeta}_{t}\right)$ (3.14)
$\quad\qquad-2\,\phi\,\left(\rho\,Q^{B,\nu}_{t}+(1-\rho)\,Q^{B,\zeta}_{t}\right)\,\Big{(}\rho\,\nu_{t}+(1-\rho)\,\zeta_{t}-\eta_{t}-\xi_{t}\Big{)}-r^{B}\,\left(\rho\,Q^{B,\nu}_{t}+(1-\rho)\,Q^{B,\zeta}_{t}\right)^{2}\bigg{\}}\,\mathrm{d}t\Bigg{]}$
$=\rho\,J^{B}(\nu,\eta)+\rho\,(1-\rho)\,\mathbb{E}\left[\int_{0}^{T}\left\{a\,\nu^{2}_{t}+\mathfrak{p}\,Q^{B,\nu}_{t}\,Y^{\nu}_{t}-\mathfrak{h}\,Q^{B,\nu}_{t}\,\nu_{t}+2\,\phi\,Q^{B,\nu}_{t}\,\nu_{t}+r^{B}\,\left(Q^{B,\nu}_{t}\right)^{2}\right\}\mathrm{d}t\right]$
$\quad+(1-\rho)\,J^{B}(\zeta,\eta)+\rho\,(1-\rho)\,\mathbb{E}\left[\int_{0}^{T}\left\{a\,\zeta^{2}_{t}-\mathfrak{h}\,Q^{B,\zeta}_{t}\,\zeta_{t}+\mathfrak{p}\,Q^{B,\zeta}_{t}\,Y^{\zeta}_{t}+2\,\phi\,Q^{B,\zeta}_{t}\,\zeta_{t}+r^{B}\,\left(Q^{B,\zeta}_{t}\right)^{2}\right\}\mathrm{d}t\right]$
$\quad-\rho\,(1-\rho)\,\mathbb{E}\Bigg{[}\int_{0}^{T}\bigg{\{}2\,a\,\nu_{t}\,\zeta_{t}-\mathfrak{h}\,\left(Q^{B,\zeta}_{t}\,\nu_{t}+Q^{B,\nu}_{t}\,\zeta_{t}\right)+\mathfrak{p}\,\left(Q^{B,\zeta}_{t}\,Y^{\nu}_{t}+Q^{B,\nu}_{t}\,Y^{\zeta}_{t}\right)$ (3.15)
$\qquad\hskip 140.00021pt+2\,\phi\,\left(Q^{B,\nu}_{t}\,\zeta_{t}+Q^{B,\zeta}_{t}\,\nu_{t}\right)+2\,r^{B}\,Q^{B,\nu}_{t}\,Q^{B,\zeta}_{t}\bigg{\}}\,\mathrm{d}t\Bigg{]}$
$=\rho\,J^{B}(\nu,\eta)+(1-\rho)\,J^{B}(\zeta,\eta)$
$\qquad+\rho\,(1-\rho)\,\mathbb{E}\Bigg{[}\int_{0}^{T}\bigg{\{}a\left(\nu_{t}-\zeta_{t}\right)^{2}+(2\,\phi-\mathfrak{h})\,\left(Q^{B,\nu}_{t}-Q^{B,\zeta}_{t}\right)\left(\nu_{t}-\zeta_{t}\right)$ (3.16)
$\hskip 110.00017pt+\mathfrak{p}\,\left(Q^{B,\nu}_{t}-Q^{B,\zeta}_{t}\right)\left(Y^{\nu}_{t}-Y^{\zeta}_{t}\right)+r^{B}\left(Q^{B,\nu}_{t}-Q^{B,\zeta}_{t}\right)^{2}\bigg{\}}\,\mathrm{d}t\Bigg{]}\,,$ (3.17)
because $\mu(A)>0$ and $\mathbb{E}\left[\int_{0}^{T}\left(\nu_{t}-\zeta_{t}\right)^{2}\mathrm{d}t\right]>0$. Further, as in (3.9),
$\mathbb{E}\left[\textstyle\int_{0}^{T}\left(Q^{B,\nu}_{t}-Q^{B,\zeta}_{t}\right)\left(\nu_{t}-\zeta_{t}\right)\mathrm{d}t\right]\allowbreak=\tfrac{1}{2}\,\mathbb{E}\left[\left(\textstyle\int_{0}^{T}(\nu_{t}-\zeta_{t})\,\mathrm{d}t\right)^{2}\right]\geq 0.$
Moreover,
$\Bigg{|}\mathbb{E}\left[\textstyle\int_{0}^{T}\left(Q^{B,\nu}_{t}-Q^{B,\zeta}_{t}\right)\left(Y^{\nu}_{t}-Y^{\zeta}_{t}\right)\mathrm{d}t\right]\Bigg{|}$ $\leq\mathbb{E}\left[\textstyle\int_{0}^{T}\left(\textstyle\int_{0}^{t}|\nu_{s}-\zeta_{s}|\,\mathrm{d}s\right)\left(\mathfrak{h}\,\textstyle\int_{0}^{t}e^{-\mathfrak{p}\,(t-u)}|\nu_{u}-\zeta_{u}|\mathrm{d}u\right)\mathrm{d}t\right]$
$\leq T\,\mathfrak{h}\,\mathbb{E}\left[\textstyle\int_{0}^{T}\left(\nu_{s}-\zeta_{s}\right)^{2}\,\mathrm{d}s\right]\,.$ (3.18)
With these inequalities and (3.17), it follows that
$J^{B}(\rho\,\nu+(1-\rho)\,\zeta,\eta)>\rho\,J^{B}(\nu,\eta)+(1-\rho)\,J^{B}(\zeta,\eta)\,,$ (3.19)
because $2\,\phi-\mathfrak{h}\geq 0$ and $a>\mathfrak{p}\,\mathfrak{h}\,T$, which concludes the proof. ∎
Armed with strict the concavity of both $J^{I}$ and $J^{B}$ (in the relevant variables), we study the Gâteaux derivatives of both functionals.
Proposition 3.3 (Gâteaux derivative of informed trader’s functional).
Let $\nu\in{\mathbb{H}^{2}}$ and $\eta,\eta^{\prime}\in{\mathbb{H}^{2}}$. The Gâteaux derivative of $J^{I}$ at $\eta$ in the direction of $\mathfrak{n}:=(\eta^{\prime}-\eta)$ is given by
$\langle\mathcal{D}J^{I}(\nu,\eta),\mathfrak{n}\rangle=\mathbb{E}\left[\int_{0}^{T}\mathfrak{n}_{t}\left\{-2\,b\,\eta_{t}-2\,\psi\,Q^{I}_{t}+\int_{t}^{T}\left(\alpha_{u}+\mathfrak{h}\,\nu_{t}-\mathfrak{p}\,Y_{t}-2\,\psi\,\eta_{u}-2\,r^{I}\,Q^{I}_{u}\right)\mathrm{d}u\right\}\mathrm{d}\right]\,.$ (3.20)
Proof.
Let $\eta,\eta^{\prime}\in{\mathbb{H}^{2}}$ and $\mathfrak{n}:=(\eta^{\prime}-\eta)$, let $\nu\in{\mathbb{H}^{2}}$, and let $\varepsilon>0$. Next, for $t\in{\mathfrak{T}}$ we have
$Q^{I,\eta+\varepsilon\,\mathfrak{n}}_{t}=Q^{I,\eta}_{t}+\varepsilon\,\int_{0}^{t}\mathfrak{n}_{u}\,\mathrm{d}u$ (3.21)
and
$J^{I}(\nu,\eta+\varepsilon\,\mathfrak{n})-J^{I}(\nu,\eta)$ $=\mathbb{E}\Bigg{[}\int_{0}^{T}\bigg{\{}-2\,b\,\varepsilon\,\eta_{t}\,\mathfrak{n}_{t}-b\,\varepsilon^{2}\,\mathfrak{n}^{2}_{t}+\varepsilon\,\int_{0}^{t}\mathfrak{n}_{u}\,\mathrm{d}u\,(\alpha_{t}+\mathfrak{h}\,\nu_{t}-\mathfrak{p}\,Y_{t})$
$\qquad\qquad-2\,\psi\,\left(Q^{I,\eta}_{t}\,\varepsilon\,\mathfrak{n}_{t}+\varepsilon\,\int_{0}^{t}\mathfrak{n}_{u}\,\mathrm{d}u\,\,\eta_{t}+\varepsilon^{2}\,\mathfrak{n}_{t}\,\int_{0}^{t}\mathfrak{n}_{u}\,\mathrm{d}u\right)$
$\qquad\qquad\qquad-r^{I}\,\left(2\,\varepsilon\,Q^{I,\eta}_{t}\,\int_{0}^{t}\mathfrak{n}_{u}\,\mathrm{d}u+\varepsilon^{2}\,\left(\int_{0}^{t}\mathfrak{n}_{u}\,\mathrm{d}u\right)^{2}\right)\bigg{\}}\,\mathrm{d}t\Bigg{]}$
$\quad=\varepsilon\,\mathbb{E}\Bigg{[}\int_{0}^{T}\bigg{\{}-2\,b\,\eta_{t}\,\mathfrak{n}_{t}+\int_{0}^{t}\mathfrak{n}_{u}\,\mathrm{d}u\,(\alpha_{t}+\mathfrak{h}\,\nu_{t}-\mathfrak{p}\,Y_{t})$
$\qquad\qquad-2\,\psi\,\left(Q^{I,\eta}_{t}\,\mathfrak{n}_{t}+\int_{0}^{t}\mathfrak{n}_{u}\,\mathrm{d}u\,\eta_{t}\right)-2\,r^{I}\,Q^{I,\eta}_{t}\,\int_{0}^{t}\mathfrak{n}_{u}\,\mathrm{d}u\bigg{\}}\,\mathrm{d}t\Bigg{]}+\mathcal{O}\left(\varepsilon^{2}\right)\,.$
The coefficient for $\varepsilon^{2}$ is finite by Lemma 2.1. It then follows that
$\langle\mathcal{D}\,J^{I}(\nu,\eta),\mathfrak{n}\rangle$ $=\mathbb{E}\Bigg{[}\int_{0}^{T}\bigg{\{}-2\,b\,\eta_{t}\,\mathfrak{n}_{t}+\int_{0}^{t}\mathfrak{n}_{u}\,\mathrm{d}u\,\left(\alpha_{t}+\mathfrak{h}\,\nu_{t}-\mathfrak{p}\,Y_{t}-2\,\psi\,\eta_{t}-2\,r^{I}\,Q^{I,\eta}_{t}\right)-2\,\psi\,Q^{I,\eta}_{t}\,\mathfrak{n}_{t}\bigg{\}}\,\mathrm{d}t\Bigg{]}$
$=\mathbb{E}\Bigg{[}\int_{0}^{T}\mathfrak{n}_{t}\,\bigg{\{}-2\,b\,\eta_{t}-2\,\psi\,Q^{I,\eta}_{t}+\int_{t}^{T}\left(\alpha_{u}+\mathfrak{h}\,\nu_{t}-\mathfrak{p}\,Y_{t}-2\,\psi\,\eta_{u}-2\,r^{I}\,Q^{I,\eta}_{u}\right)\mathrm{d}u\bigg{\}}\,\mathrm{d}t\Bigg{]}\,,$ (3.22)
where the last line is a consequence of Fubini’s theorem. ∎
Proposition 3.4 (Gâteaux derivative of broker’s functional).
Let $\nu,\nu^{\prime}\in{\mathbb{H}^{2}}$ and $\eta\in{\mathbb{H}^{2}}$. The Gâteaux derivative of $J^{B}$ at $\nu$ in the direction of $\mathfrak{v}:=(\nu^{\prime}-\nu)$ is given by
$\langle\mathcal{D}\,J^{B}(\nu,\eta),\mathfrak{v}\rangle=\mathbb{E}\Bigg{[}$ $\int_{0}^{T}\mathfrak{v}_{t}\,\bigg{\{}-2\,a\,\nu_{t}-(2\,\phi-\mathfrak{h})\,Q^{B,\nu}_{t}-\mathfrak{p}\,\mathfrak{h}\,\int_{t}^{T}Q^{B,\nu}_{u}\,e^{-\mathfrak{p}\,(u-t)}\,\mathrm{d}u$ (3.23)
$\quad+\int_{t}^{T}\left(\alpha_{u}-(2\,\phi-\mathfrak{h})\,\nu_{u}-\mathfrak{p}\,Y_{u}+2\,\phi\,\left(\eta_{u}+\xi_{u}\right)-2\,r^{B}\,Q^{B,\nu}_{u}\right)\mathrm{d}u\bigg{\}}\,\mathrm{d}t\Bigg{]}\,.$
Proof.
Let $\eta\in{\mathbb{H}^{2}}$, let $\nu,\nu^{\prime}\in{\mathbb{H}^{2}}$ and $\mathfrak{v}:=(\nu^{\prime}-\nu)$, and let $\varepsilon>0$. Next, for $t\in{\mathfrak{T}}$ we have
$Q^{B,\nu+\varepsilon\,\mathfrak{v}}_{t}=Q^{B,\nu}_{t}+\varepsilon\,\int_{0}^{t}\mathfrak{v}_{u}\,\mathrm{d}u\,,$ (3.24)
and
$Y^{\nu+\varepsilon\,\mathfrak{v}}_{t}=Y^{\nu}_{t}+\varepsilon\,\mathfrak{h}\,\int_{0}^{t}e^{-\mathfrak{p}\,(t-u)}\,\mathfrak{v}_{u}\,\mathrm{d}u\,,$ (3.25)
therefore,
$J^{B}(\nu+\varepsilon\,\mathfrak{v},\eta)-J^{B}(\nu,\eta)$ $=\varepsilon\,\mathbb{E}\Bigg{[}\int_{0}^{T}\Bigg{\{}-2\,a\,\nu_{t}\,\mathfrak{v}_{t}+\alpha_{t}\,\int_{0}^{t}\mathfrak{v}_{u}\mathrm{d}u$
$\qquad\qquad-(2\,\phi-\mathfrak{h})\,\left(Q^{B,\nu}_{t}\,\mathfrak{v}_{t}+\nu_{t}\,\int_{0}^{t}\mathfrak{v}_{u}\,\mathrm{d}u\right)$
$\qquad\qquad-\mathfrak{p}\,\mathfrak{h}\,Q^{B,\nu}_{t}\,\int_{0}^{t}e^{-\mathfrak{p}\,(t-s)}\,\mathfrak{v}_{s}\,\mathrm{d}s-\mathfrak{p}\,Y_{t}\,\int_{0}^{t}\mathfrak{v}_{s}\,\mathrm{d}s$
$\qquad\qquad+2\,\phi\,\left(\eta_{t}+\xi_{t}\right)\,\int_{0}^{t}\mathfrak{v}_{u}\,\mathrm{d}u-2\,r^{B}\,Q^{B,\nu}_{t}\,\int_{0}^{t}\mathfrak{v}_{u}\,\mathrm{d}u\bigg{\}}\,\mathrm{d}t\Bigg{]}+\mathcal{O}\left(\varepsilon^{2}\right)\,.$
After applying Fubini’s theorem, it follows that
$\langle\mathcal{D}\,J^{B}(\nu,\eta),\mathfrak{v}\rangle$ $=\mathbb{E}\Bigg{[}\int_{0}^{T}\mathfrak{v}_{t}\,\bigg{\{}-2\,a\,\nu_{t}-(2\,\phi-\mathfrak{h})\,Q^{B,\nu}_{t}-\mathfrak{p}\,\mathfrak{h}\,\int_{t}^{T}Q^{B,\nu}_{u}\,e^{-\mathfrak{p}\,(u-t)}\,\mathrm{d}u$ (3.26)
$\hskip 65.00009pt+\int_{t}^{T}\left(\alpha_{u}-(2\,\phi-\mathfrak{h})\,\nu_{u}-\mathfrak{p}\,Y_{u}+2\,\phi\,\left(\eta_{u}+\xi_{u}\right)-2\,r^{B}\,Q^{B,\nu}_{u}\right)\mathrm{d}u\bigg{\}}\,\mathrm{d}t\Bigg{]}\,,$
which concludes the proof. ∎
The following theorem characterises the best response of the informed trader as the solution to an FBSDE.
Theorem 3.5.
Let $\nu\in{\mathbb{H}^{2}}$. A control $\eta^{*}\in{\mathbb{H}^{2}}$ maximises the functional $\eta\to J^{I}(\nu,\eta)$ if and only if
$\begin{cases}&\mathrm{d}\eta^{*}_{t}=-\frac{1}{2\,b}\left(\alpha_{t}-2\,r^{I}\,Q^{I,\eta^{*}}_{t}+\mathfrak{h}\,\nu_{t}-\mathfrak{p}\,Y_{t}\right)\mathrm{d}t+\mathrm{d}M^{I}_{t}\,,\\[5.0pt] &\eta^{*}_{T}=-\frac{\psi}{b}\,Q^{I,\eta^{*}}_{T}\,,\end{cases}$ (3.27)
for a suitable ${\mathcal{F}}$-martingale $(M^{I}_{t})_{t\in{\mathfrak{T}}}$.
Proof.
Let $\nu\in{\mathbb{H}^{2}}$. By Proposition 3.1, $\eta\to J^{I}(\nu,\eta)$ is a strictly concave functional and by Proposition 3.3 it follows from Ekeland and Temam, (1999) (Proposition 2.1 in Chapter 2) that $\eta^{*}\in{\mathbb{H}^{2}}$ maximises the functional $J^{I}$ if and only if
$\langle\mathcal{D}\,J^{I}(\nu,\eta^{*}),(\eta^{\prime}-\eta^{*})\rangle=0\,,\quad\forall\;\eta^{\prime}\in{\mathbb{H}^{2}}\,.$ (3.28)
Next, we show that this happens if and only if $\eta^{*}$ satisfies the FBSDE in (LABEL:eq:_FBSDE_1_Informed). Necessity: Let $\eta^{*}\in{\mathbb{H}^{2}}$ maximise $\eta\to J^{I}(\nu,\eta)$, so (3.28) holds. Then, for $\eta^{\prime}\in{\mathbb{H}^{2}}$ arbitrary, from Proposition 3.3 we have
$\mathbb{E}\Bigg{[}\int_{0}^{T}\mathfrak{n}_{t}\,\bigg{\{}-2\,b\,\eta^{*}_{t}-2\,\psi\,Q^{I,\eta^{*}}_{t}+\mathbb{E}\bigg{[}\int_{t}^{T}\chi_{u}^{\eta^{*}}\mathrm{d}u\,\bigg{|}\,{\mathcal{F}}_{t}\bigg{]}\bigg{\}}\,\mathrm{d}t\Bigg{]}=0\,,$ (3.29)
where $\mathfrak{n}:=(\eta^{\prime}-\eta^{*})$ and where $\chi_{u}^{\eta^{*}}:=(\alpha_{u}+\mathfrak{h}\,\nu_{u}-\mathfrak{p}\,{Y}_{u}-2\,\psi\,\eta^{*}_{u}-2\,r^{I}\,Q^{I,\eta^{*}}_{u})$. This is only possible if
$2\,b\,\eta^{*}_{t}$ $=-2\,\psi\,Q^{I,\eta^{*}}_{t}+\mathbb{E}\bigg{[}\int_{t}^{T}\chi_{u}^{\eta^{*}}\mathrm{d}u\,\bigg{|}\,{\mathcal{F}}_{t}\bigg{]}$ (3.30)
up to $\mathbb{P}\otimes\mathbb{L}$ null sets. Next, define the ${\mathcal{F}}$-martingale $(M^{I}_{t})_{t\in{\mathfrak{T}}}$ by
$M^{I}_{t}:=\mathbb{E}\bigg{[}\int_{0}^{T}\chi_{u}^{\eta^{*}}\mathrm{d}u\,\bigg{|}\,{\mathcal{F}}_{t}\bigg{]}\,,$ (3.31)
for $t\in{\mathfrak{T}}$. As $\alpha,Y,\nu,\eta^{*},Q^{I,\eta^{*}}$ are square-integrable and $\chi_{u}^{\eta^{*}}$ is a linear combination of square-integrable adapted processes, we have that
$\mathbb{E}\left[\left|\int_{0}^{T}\chi_{u}^{\eta^{*}}\mathrm{d}u\right|\;\right]\leq\left(\mathbb{E}\left[\left(\int_{0}^{T}\chi_{u}^{\eta^{*}}\,\mathrm{d}u\right)^{2}\right]\right)^{\tfrac{1}{2}}\leq\left(T\;\mathbb{E}\left[\int_{0}^{T}(\chi_{u}^{\eta^{*}})^{2}\,\mathrm{d}u\right]\right)^{\tfrac{1}{2}}<+\infty\,,$ (3.32)
hence, $(M^{I}_{t})_{t\in{\mathfrak{T}}}$ is a martingale. Thus we obtain the representation
$2\,b\,\eta^{*}_{t}$ $=-2\,\psi\,Q^{I,\eta^{*}}_{t}-\int_{0}^{t}\chi_{u}^{\eta^{*}}\mathrm{d}u+M^{I}_{t}\,,$ (3.33)
and it follows that
$2\,b\,\mathrm{d}\eta^{*}_{t}$ $=-2\,\psi\,\mathrm{d}Q^{I,\eta^{*}}_{t}-\chi_{u}^{\eta^{*}}\mathrm{d}t+\mathrm{d}M^{I}_{t}$ (3.34)
$=-\left(\alpha_{t}+\mathfrak{h}\,\nu_{t}-\mathfrak{p}\,{Y}_{t}-2\,r^{I}\,Q^{I,\eta^{*}}_{t}\right)\mathrm{d}t+\mathrm{d}M^{I}_{t}\,.$ (3.35)
Sufficiency: Let $\eta^{*}\in{\mathbb{H}^{2}}$ satisfy (LABEL:eq:_FBSDE_1_Informed). The unique strong solution to this linear FBSDE is
$\eta^{*}_{t}$ $=-\frac{\psi}{b}\,Q^{I,\eta^{*}}_{t}+\frac{1}{2\,b}\mathbb{E}\bigg{[}\int_{t}^{T}\chi_{u}^{\eta^{*}}\,\mathrm{d}u\,\bigg{|}\,{\mathcal{F}}_{t}\bigg{]}\,.$ (3.36)
Then, for $\mathfrak{n}:=(\eta^{\prime}-\eta^{*})$ with $\eta^{\prime}\in{\mathbb{H}^{2}}$ arbitrary,
$\langle\mathcal{D}\,J^{I}(\nu,\eta^{*}),\mathfrak{n}\rangle$ $=\mathbb{E}\left[\int_{0}^{T}\mathfrak{n}_{t}\left\{-2\,b\,\eta^{*}_{t}-2\,\psi\,Q^{I,\eta^{*}}_{t}+\int_{t}^{T}\chi_{u}^{\eta^{*}}\mathrm{d}u\right\}\mathrm{d}t\right]$
$=\mathbb{E}\Bigg{[}\int_{0}^{T}\mathfrak{n}_{t}\bigg{\{}-\mathbb{E}\bigg{[}\int_{t}^{T}\chi_{u}^{\eta^{*}}\mathrm{d}u\,\bigg{|}\,{\mathcal{F}}_{t}\bigg{]}+\int_{t}^{T}\chi_{u}^{\eta^{*}}\mathrm{d}u\bigg{\}}\,\mathrm{d}t\Bigg{]}$
$=\mathbb{E}\Bigg{[}\int_{0}^{T}\mathfrak{n}_{t}\bigg{\{}-\mathbb{E}\bigg{[}\int_{0}^{T}\chi_{u}^{\eta^{*}}\mathrm{d}u\,\bigg{|}\,{\mathcal{F}}_{t}\bigg{]}+\int_{0}^{T}\chi_{u}^{\eta^{*}}\mathrm{d}u\bigg{\}}\,\mathrm{d}t\Bigg{]}$
$=\mathbb{E}\Bigg{[}\int_{0}^{T}\mathfrak{n}_{t}\bigg{\{}-M^{I}_{t}+M^{I}_{T}\bigg{\}}\,\mathrm{d}t\Bigg{]}=\mathbb{E}\Bigg{[}\int_{0}^{T}\mathfrak{n}_{t}\bigg{\{}\mathbb{E}\bigg{[}-M^{I}_{t}+M^{I}_{T}\,\Big{|}\,{\mathcal{F}}_{t}\bigg{]}\bigg{\}}\,\mathrm{d}t\Bigg{]}=0\,,$
Thus, $\eta^{*}$ is optimal because
$\langle\mathcal{D}\,J^{I}(\nu,\eta^{*}),(\eta^{\prime}-\eta^{*})\rangle=0\,,\qquad\forall\eta^{\prime}\in{\mathbb{H}^{2}}\qquad\iff\quad\eta^{*}=\operatorname*{arg\,max}_{\eta\in{\mathbb{H}^{2}}}J^{I}(\eta,\nu^{*})\,.$ (3.37)
Similar to the theorem above, the theorem below characterises the best response of the broker as the solution to an FBSDE.
Theorem 3.6.
Let $\eta\in{\mathbb{H}^{2}}$. A control $\nu^{*}\in{\mathbb{H}^{2}}$ maximises the functional $\nu\to J^{B}(\nu,\eta)$ if and only if
$\begin{cases}&\mathrm{d}\nu^{*}_{t}=-\frac{1}{2\,a}\left(\alpha_{t}-\left(2\,r^{B}+\mathfrak{p}\,\mathfrak{h}\right)Q^{B,\nu^{*}}_{t}-\mathfrak{p}\,Y_{t}+\mathfrak{h}\,\left(\eta_{t}+\xi_{t}\right)+\mathfrak{p}^{2}\,\mathfrak{h}\,Z_{t}\right)\mathrm{d}t+\frac{1}{2\,a}\mathrm{d}\tilde{M}^{B}_{t}-\frac{\mathfrak{p}\,\mathfrak{h}}{2\,a}e^{\mathfrak{p}\,t}\,\mathrm{d}\tilde{N}^{B}_{t},\\[5.0pt] &\mathrm{d}Z_{t}=\left(\mathfrak{p}\,Z_{t}-Q^{B,\nu^{*}}_{t}\right)\mathrm{d}t+e^{\mathfrak{p}\,t}\mathrm{d}\tilde{N}^{B}_{t}\,,\\[5.0pt] &Z_{T}=0\,,\\[5.0pt] &\nu^{*}_{T}=-\frac{2\,\phi-\mathfrak{h}}{2\,a}\,Q^{B,\nu^{*}}_{T}\,,\end{cases}$ (3.38)
for suitable ${\mathcal{F}}$-martingales $(\tilde{N}^{B}_{t})_{t\in{\mathfrak{T}}}$ and $(\tilde{M}^{B}_{t})_{t\in{\mathfrak{T}}}$.
Proof.
The proof is similar to that of Theorem 3.5. Let $\eta\in{\mathbb{H}^{2}}$. By Proposition 3.2, $\nu\to J^{B}(\nu,\eta)$ is a strictly concave functional and as before, by Proposition 3.4, it follows that $\nu^{*}\in{\mathbb{H}^{2}}$ maximises the functional $J^{B}$ if and only if
$\langle\mathcal{D}\,J^{B}(\nu^{*},\eta),(\nu^{\prime}-\nu^{*})\rangle=0\,,\quad\forall\nu^{\prime}\in{\mathbb{H}^{2}}\,.$ (3.39)
Next, we show that this happens if and only if $\nu^{*}$ satisfies the FBSDE in (3.38). Necessity: Let $\nu^{*}\in{\mathbb{H}^{2}}$ maximise $\nu\to J^{B}(\nu,\eta)$, then, it follows that (3.39) holds. Take $\nu^{\prime}\in{\mathbb{H}^{2}}$ arbitrary, then from Proposition 3.4, we have
00 0 $=\mathbb{E}\Bigg{[}\int_{0}^{T}\mathfrak{v}_{t}\,\bigg{\{}-2\,a\,\nu^{*}_{t}-(2\,\phi-\mathfrak{h})\,Q^{B,\nu^{*}}_{t}-\mathfrak{p}\,\mathfrak{h}\,\int_{t}^{T}Q^{B,\nu^{*}}_{u}\,e^{-\mathfrak{p}\,(u-t)}\,\mathrm{d}u+\int_{t}^{T}\chi_{u}^{\nu^{*}}\mathrm{d}u\bigg{\}}\,\mathrm{d}t\Bigg{]}\,,$
where $\mathfrak{v}:=(\nu^{\prime}-\nu^{*})$ and $\chi_{u}^{\nu^{*}}:=\left(\alpha_{u}-(2\,\phi-\mathfrak{h})\,\nu^{*}_{u}-\mathfrak{p}\,Y_{u}+2\,\phi\,\left(\eta_{u}+\xi_{u}\right)-2\,r^{B}\,Q^{B,\nu^{*}}_{u}\right)$. This is only possible if
$2\,a\,\nu^{*}_{t}$ $=-(2\,\phi-\mathfrak{h})\,Q^{B,\nu^{*}}_{t}-\mathfrak{p}\,\mathfrak{h}\,e^{\mathfrak{p}\,t}\,\mathbb{E}\left[\int_{t}^{T}Q^{B,\nu^{*}}_{u}\,e^{-\mathfrak{p}\,u}\,\mathrm{d}u\,\mid\,{\mathcal{F}}_{t}\right]+\mathbb{E}\left[\int_{t}^{T}\chi_{u}^{\nu^{*}}\mathrm{d}u\,\mid\,{\mathcal{F}}_{t}\right]$
up to $\mathbb{P}\otimes\mathbb{L}$ null sets. Hence, define the ${\mathcal{F}}$-martingales $(\tilde{N}^{B}_{t})_{t\in{\mathfrak{T}}}$ and $(\tilde{M}^{B}_{t})_{t\in{\mathfrak{T}}}$ s.t.,
$\tilde{N}^{B}_{t}:=\mathbb{E}\left[\int_{0}^{T}Q^{B,\nu^{*}}_{u}\,e^{-\mathfrak{p}\,u}\,\mathrm{d}u\,\mid\,{\mathcal{F}}_{t}\right]\qquad\text{and}\qquad\tilde{M}^{B}_{t}:=\mathbb{E}\left[\int_{0}^{T}\chi_{u}^{\nu^{*}}\mathrm{d}u\,\mid\,{\mathcal{F}}_{t}\right]\,,$ (3.40)
for $t\in{\mathfrak{T}}$. Similar to the proof that $M^{I}$ is a martingale in Theorem 3.5, $\tilde{N}^{B}$ and $\tilde{M}^{B}$ are martingales because $Q^{B,\nu^{*}}$ and $\chi_{u}^{\nu^{*}}$ are square-integrable. In particular, we have that
$\mathbb{E}\left[\left|\int_{0}^{T}Q^{B,\nu^{*}}_{u}\,e^{-\mathfrak{p}\,u}\mathrm{d}u\right|\right]<+\infty\qquad\text{and}\qquad\mathbb{E}\left[\left|\int_{0}^{T}\chi_{u}^{\nu^{*}}\mathrm{d}u\right|\right]<+\infty\,.$ (3.41)
Hence, we obtain the representation
$2\,a\,\nu^{*}_{t}$ $=-(2\,\phi-\mathfrak{h})\,Q^{B,\nu^{*}}_{t}-\mathfrak{p}\,\mathfrak{h}\,e^{\mathfrak{p}\,t}\,\left(\tilde{N}^{B}_{t}-\int_{0}^{t}Q^{B,\nu^{*}}_{u}\,e^{-\mathfrak{p}\,u}\,\mathrm{d}u\right)-\int_{0}^{t}\chi_{u}^{\nu^{*}}\mathrm{d}u+\tilde{M}^{B}_{t}\,.$ (3.42)
Next, define the auxiliary process
$Z_{t}=e^{\mathfrak{p}\,t}\,\left(\tilde{N}^{B}_{t}-\int_{0}^{t}Q^{B,\nu^{*}}_{u}\,e^{-\mathfrak{p}\,u}\,\mathrm{d}u\right)\,,$ (3.43)
which satisfies the BSDE
$\mathrm{d}Z_{t}=\left(\mathfrak{p}\,Z_{t}-Q^{B,\nu^{*}}_{t}\right)\mathrm{d}t+e^{\mathfrak{p}\,t}\mathrm{d}\tilde{N}^{B}_{t}\,,\qquad Z_{T}=0\,.$ (3.44)
Thus,
$2\,a\,\mathrm{d}\nu^{*}_{t}$ $=-(2\,\phi-\mathfrak{h})\,\left(\nu^{*}_{t}-\eta_{t}-\xi_{t}\right)\mathrm{d}t-\mathfrak{p}\,\mathfrak{h}\,\left[\left(\mathfrak{p}\,Z_{t}-Q^{B,\nu^{*}}_{t}\right)\mathrm{d}t+e^{\mathfrak{p}\,t}\mathrm{d}\tilde{N}^{B}_{t}\right]-\chi_{t}^{\nu^{*}}\mathrm{d}t+\mathrm{d}\tilde{M}^{B}_{t}\,,$ (3.45)
which simplifies to
$\mathrm{d}\nu^{*}_{t}$ $=-\frac{1}{2\,a}\left(\alpha_{t}-\left(2\,r^{B}+\mathfrak{p}\,\mathfrak{h}\right)Q^{B,\nu^{*}}_{t}-\mathfrak{p}\,Y_{t}+\mathfrak{h}\,\left(\eta_{t}+\xi_{t}\right)+\mathfrak{p}^{2}\,\mathfrak{h}\,Z_{t}\right)\mathrm{d}t+\frac{1}{2\,a}\mathrm{d}\tilde{M}^{B}_{t}-\frac{\mathfrak{p}\,\mathfrak{h}}{2\,a}e^{\mathfrak{p}\,t}\mathrm{d}\tilde{N}^{B}_{t}\,,$ (3.46)
and together with (3.1) evaluated at $T$ show that the FBSDE system is satisfied. Sufficiency: Let $\nu^{*}\in{\mathbb{H}^{2}}$ satisfy (3.38). The unique strong solution to this linear FBSDE is
$\nu^{*}_{t}$ $=-\frac{(2\,\phi-\mathfrak{h})}{2\,a}\,Q^{B,\nu^{*}}_{t}-\frac{\mathfrak{p}\,\mathfrak{h}\,e^{\mathfrak{p}\,t}}{2\,a}\,\mathbb{E}\left[\int_{t}^{T}Q^{B,\nu^{*}}_{u}\,e^{-\mathfrak{p}\,u}\,\mathrm{d}u\,\mid\,{\mathcal{F}}_{t}\right]+\frac{1}{2\,a}\mathbb{E}\left[\int_{t}^{T}\chi_{u}^{\nu^{*}}\mathrm{d}u\,\mid\,{\mathcal{F}}_{t}\right]\,,$
which, using Proposition 3.4 yields,
$\langle\mathcal{D}\,J^{B}(\nu^{*},\eta),\mathfrak{v}\rangle$ $=\mathbb{E}\Bigg{[}\int_{0}^{T}\mathfrak{v}_{t}\bigg{\{}\mathfrak{p}\,\mathfrak{h}\,e^{\mathfrak{p}\,t}\left(\tilde{N}^{B}_{t}-\tilde{N}^{B}_{T}\right)-\tilde{M}^{B}_{t}+\tilde{M}^{B}_{T}\bigg{\}}\,\mathrm{d}t\Bigg{]}$
$=\mathbb{E}\Bigg{[}\int_{0}^{T}\mathfrak{v}_{t}\bigg{\{}\mathfrak{p}\,\mathfrak{h}\,e^{\mathfrak{p}\,t}\mathbb{E}\bigg{[}\tilde{N}^{B}_{t}-\tilde{N}^{B}_{T}\,\Big{|}\,{\mathcal{F}}_{t}\bigg{]}+\mathbb{E}\bigg{[}\tilde{M}^{B}_{T}-\tilde{M}^{B}_{t}\,\Big{|}\,{\mathcal{F}}_{t}\bigg{]}\bigg{\}}\,\mathrm{d}t\Bigg{]}=0\,,$
for all $\nu^{\prime}\in{\mathbb{H}^{2}}$ and $\mathfrak{v}:=\nu^{\prime}-\nu^{*}$. Thus, it follows that $\nu^{*}$ is optimal. ∎
The next theorem shows that if there is a solution to a coupled system of FBSDEs, then, the solution to the system satisfies Definition 2.2, and thus, it is a Nash equilibrium of the stochastic game.
Theorem 3.7.
If there is $\nu^{*}\in{\mathbb{H}^{2}}$ and $\eta^{*}\in{\mathbb{H}^{2}}$ that satisfy the system of FBSDEs
$\mathrm{d}\nu^{*}_{t}=-\frac{1}{2\,a}\left(\alpha_{t}-\left(2\,r^{B}+\mathfrak{p}\,\mathfrak{h}\right)Q^{B,\nu^{*}}_{t}-\mathfrak{p}\,Y_{t}+\mathfrak{h}\,\left(\eta^{*}_{t}+\xi_{t}\right)+\mathfrak{p}^{2}\,\mathfrak{h}\,Z_{t}\right)\mathrm{d}t$
$\hskip 227.62204pt+\frac{1}{2\,a}\mathrm{d}\tilde{M}^{B}_{t}-\frac{\mathfrak{p}\,\mathfrak{h}}{2\,a}e^{\mathfrak{p}\,t}\,\mathrm{d}\tilde{N}^{B}_{t}\,,$ (3.47a)
$\nu^{*}_{T}=-\frac{\varphi}{a}\,Q^{B,\nu^{*}}_{T}\,,$ (3.47b)
$\mathrm{d}\eta^{*}_{t}=-\frac{1}{2\,b}\left(\alpha_{t}-2\,r^{I}\,Q^{I,\eta^{*}}_{t}+\mathfrak{h}\,\nu^{*}_{t}-\mathfrak{p}\,Y_{t}\right)\mathrm{d}t+\mathrm{d}M^{I}_{t}\,,$ (3.47c)
$\eta^{*}_{T}=-\frac{\psi}{b}\,Q^{I,\eta^{*}}_{T}\,,$ (3.47d)
$\mathrm{d}Y_{t}=(\mathfrak{h}\,\nu^{*}_{t}-\mathfrak{p}\,Y_{t})\,\mathrm{d}t\,,\quad Y_{0}\in\mathbb{R}\,,$ (3.47e)
$\mathrm{d}Z_{t}=\left(\mathfrak{p}\,Z_{t}-Q^{B,\nu^{*}}_{t}\right)\mathrm{d}t+e^{\mathfrak{p}\,t}\mathrm{d}\tilde{N}^{B}_{t}\,,\qquad Z_{T}=0\,,$ (3.47f)
$\mathrm{d}Q^{B,\nu^{*}}_{t}=\left(\nu^{*}_{t}-\eta^{*}_{t}-\xi_{t}\right)\mathrm{d}t\,,\qquad Q^{B,\nu^{*}}_{0}=q^{B}_{0}\,,$ (3.47g)
$\mathrm{d}Q^{I,\eta^{*}}_{t}=\eta^{*}_{t}\,\mathrm{d}t\,,\qquad Q^{I,\eta^{*}}_{0}=q^{I}_{0}\,,$ (3.47h)
where $\varphi=\phi-\tfrac{1}{2}\mathfrak{h}$, and the processes $(\tilde{N}^{B}_{t})_{t\in{\mathfrak{T}}}$ and $(\tilde{M}^{B}_{t})_{t\in{\mathfrak{T}}}$ are ${\mathcal{F}}$-martingales, and $(M^{I}_{t})_{t\in{\mathfrak{T}}}$ is a ${\mathcal{F}}$-martingale. Then, $(\nu^{*},\eta^{*})\in{\mathbb{H}^{2}}\times{\mathbb{H}^{2}}$ is a Nash equilibrium of the trading strategies between the broker and the informed trader.
Proof.
Suppose there is $\eta^{*}\in{\mathbb{H}^{2}}$ and $\nu^{*}\in{\mathbb{H}^{2}}$ such that the system of FBSDEs (3.47) holds; then, by Theorem 3.5 it follows that
$J^{I}(\eta^{*},\nu^{*})\geq J^{I}(\kappa,\nu^{*})\,,\qquad\forall\;\kappa\in{\mathbb{H}^{2}}\,.$ (3.48)
Similarly, by Theorem 3.6, it follows that
$J^{B}(\eta^{*},\nu^{*})\geq J^{B}(\eta^{*},\zeta)\,,\qquad\forall\;\zeta\in{\mathbb{H}^{2}}\,.$ (3.49)
Thus, $(\eta^{*},\nu^{*})$ are a Nash equilibrium because they satisfy Definition 2.2. Uniqueness up to $\mathbb{P}\otimes\mathbb{L}$ null sets follows from the strict concavity of both $\eta\to J^{I}(\eta,\nu^{*})$ and $\nu\to J^{B}(\eta^{*},\nu)$. ∎
3.2 Existence and uniqueness
In this section we are concerned with the existence and uniqueness of the system of FBSDEs above that characterises the Nash equilirbium. Introduce the (vector valued) forward process $(\mathcal{X}_{t})_{t\in{\mathfrak{T}}}$, backward process $(\mathcal{Y}_{t})_{t\in{\mathfrak{T}}}$, and martingale process $(\mathcal{M}_{t})_{t\in{\mathfrak{T}}}$, where
$\mathcal{X}_{t}$ $=\begin{pmatrix}Q^{B,\nu^{*}}_{t}\\ Q^{I,\eta^{*}}_{t}\\ Y_{t}\end{pmatrix}\,,\qquad\mathcal{Y}_{t}=\begin{pmatrix}\nu^{*}_{t}\\ \eta^{*}_{t}\\ Z_{t}\end{pmatrix}\,,\qquad\text{and}\qquad\mathcal{M}_{t}=\begin{pmatrix}\tilde{N}^{B}_{t}\\ \tilde{M}^{B}_{t}\\ M^{I}_{t}\end{pmatrix}\,.$ (3.50)
The FBSDE in (3.47) can be written compactly as follows
$\mathrm{d}\mathcal{X}_{t}$ $=\left(A\,\mathcal{X}_{t}+B\,\mathcal{Y}_{t}+b_{t}\right)\,\mathrm{d}t\,,$ $\mathcal{X}_{0}$ $=\left(q^{B}_{0},q^{I}_{0},Y_{0}\right)^{\intercal},$ (3.51a)
$\mathrm{d}\mathcal{Y}_{t}$ $=\left(\hat{A}\,\mathcal{X}_{t}+\hat{B}\,\mathcal{Y}_{t}+\hat{b}_{t}\right)\,\mathrm{d}t+\sigma_{t}\,\mathrm{d}\mathcal{M}_{t}\,,$ $\mathcal{Y}_{T}$ $=G\,\mathcal{X}_{T}\,,$ (3.51b)
where $A,B,\hat{A},\hat{B},G$ are deterministic matrices in $\mathbb{R}^{3\times 3}$, the stochastic processes $b_{t},\hat{b}_{t}:{\mathfrak{T}}\times\Omega\to\mathbb{R}^{3}$ are valued in $\mathbb{R}^{3}$, and $\sigma_{t}:{\mathfrak{T}}\to\mathbb{R}^{3\times 3}$ is a deterministic function, all of which we obtain from (3.47). More precisely,
[…]
$\hat{A}=\begin{pmatrix}\frac{2\,r^{B}+\mathfrak{p}\,\mathfrak{h}}{2a}&0&\frac{\mathfrak{p}}{2\,a}\vspace{0.1cm}\\ 0&\frac{r^{I}}{b}&\frac{\mathfrak{p}}{2\,b}\vspace{0.1cm}\\ -1&0&0\end{pmatrix},\qquad\hat{B}=\begin{pmatrix}0&-\frac{\mathfrak{h}}{2\,a}&-\frac{\mathfrak{p}^{2}\,\mathfrak{h}}{2\,a}\vspace{0.1cm}\\ -\frac{\mathfrak{h}}{2\,b}&0&0\vspace{0.1cm}\\ 0&0&\mathfrak{p}\,\end{pmatrix},$ (3.53)
$G=\begin{pmatrix}-\frac{\varphi}{a}&0&0\\ 0&-\frac{\psi}{b}&0\\ 0&0&0\end{pmatrix},\qquad\hat{b}_{t}=\begin{pmatrix}-\frac{\alpha_{t}+\mathfrak{h}\,\xi_{t}}{2\,a}\vspace{0.1cm}\\ -\frac{\alpha_{t}}{2\,b}\vspace{0.1cm}\\ 0\end{pmatrix},\qquad\sigma_{t}=\begin{pmatrix}-\frac{\mathfrak{p}\,\mathfrak{h}}{2\,a}e^{\mathfrak{p}\,t}&\frac{1}{2\,a}&0\vspace{0.1cm}\\ 0&0&1\vspace{0.1cm}\\ e^{\mathfrak{p}\,t}&0&0\end{pmatrix}.$ (3.54)
If $T$ is small enough, then, existence and uniqueness of the slution to the FBSDE (3.51) is well understood in the literature; see e.g., Antonelli, (1993) or Chapter 2 in Carmona, (2016). In the Brownian case, Yong, (1999) shows that existence and uniqueness of a solution to the FBSDE system (3.51) is equivalent to the existence and uniqueness of a matrix Riccati differential equation; this result follows the ‘four-step scheme’ of Ma et al., (1994).222 See Ma and Yong, (1999); Carmona, (2016) for manuscripts covering a wide range of existence and uniqueness results for FBSDEs. Below, similar to these previous works, we take a constructive approach and show that existence and uniqueness of the solution to the FBSDE (3.51) is equivalent to the existence and uniqueness of a non-Hermitian matrix Riccati differential equation. We show that when the values of the parameters $\mathfrak{p}$ and $r^{B}$ are zero, existence and uniqueness of the solution follows without any further conditions on the remaining model parameters. Alternatively, if only $\mathfrak{p}$ is zero, existence and uniqueness follows under four additional (mild) restrictions on model parameters. In the numerical examples below we take $\mathfrak{p}=r^{B}=0$, thus, we employ the result of existence and uniqueness that we prove below with tools from the literature on non-Hermitian matrix Riccati differential equations. Finally, in the general case when $\mathfrak{p}$ is not zero and $T$ is small enough, then existence and uniqueness follow from the arguments in Antonelli, (1993).
For the sake of completeness, we show the precise constraint on $T$ to guarantee existence and uniqueness in the case of a small time horizon. [^3] As noted in page 58 in Carmona, (2016) regarding the affine FBSDE (2.50) in that page: “except for possibly in the case $C_{t}=D_{t}=0$ and $\sigma_{t}\equiv\sigma>0$, and despite the innocent-looking form of the coefficients of FBSDE (2.50), none of the existence theorems which we slaved to prove in the previous sections apply to this simple particular case”. Within the Brownian context, our FBSDE is like that in (2.50) in Carmona, (2016) but for $C_{t}=D_{t}=\sigma_{t}=0$. Let
$\mathbb{S}^{2}:=\Bigg{\{}(\gamma_{t})_{t\in{\mathfrak{T}}}\,\,|\,\,\gamma\text{ is }{\mathcal{F}}-\text{adapted, and }\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\textstyle|\eta_{t}|^{2}\right]\,<\,+\infty\Bigg{\}}\,,$ (3.55)
and define $\mathbb{S}^{2,3}=\mathbb{S}^{2}\times\mathbb{S}^{2}\times\mathbb{S}^{2}$ and let $\mathbb{K}=\mathbb{S}^{2,3}\times\mathbb{S}^{2,3}$. We endow the vector space $\mathbb{K}$ with the following norm: for $(X,Y)\in\mathbb{K}$
$\left\lVert(X,Y)\right\rVert_{\mathbb{K}}^{2}:=\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}|{X_{t}}|^{2}\right]+\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}|{Y_{t}}|^{2}\right]\,,$ (3.56)
[…]
The theorem below shows the bound that $T$ needs to satisfy so that existence and uniqueness follows from a fixed-point argument.
Theorem 3.8.
If the time horizon $T$ and $|G|$ are such that
$\max\left\{12\,|G|^{2}+T^{2}(2\,|A|^{2}+30\,|\hat{A}|^{2})\,,\,T^{2}\left(2\,|B|^{2}+30\,|\hat{B}|^{2}\right)\right\}<1\,,$ (3.58)
then, there is a unique solution to the FBSDE system in (3.47). When $|G|=0$, the inequality reduces to
[…]
Proof.
Consider the mapping $\Gamma:\mathbb{K}\to\mathbb{K}$ that maps $(\mathcal{X},\mathcal{Y})\in\mathbb{K}$ to
$\Gamma(\mathcal{X},\mathcal{Y})_{t}=\left(\mathcal{X}_{0}+\int_{0}^{t}\left(A\,\mathcal{X}_{s}+B\,\mathcal{Y}_{s}+b_{s}\right)\mathrm{d}s,\,\mathbb{E}\left[G\,\mathcal{X}_{T}-\int_{t}^{T}\left(\hat{A}\,\mathcal{X}_{s}+\hat{B}\,\mathcal{Y}_{s}+\hat{b}_{s}\right)\mathrm{d}s\,\Big{|}\,{\mathcal{F}}_{t}\right]\right)\,.$
First, we show that indeed for $(\mathcal{X},\mathcal{Y})\in\mathbb{K}$, we have that $\Gamma(\mathcal{X},\mathcal{Y})\in\mathbb{K}$. Observe that
$\left\lVert\Gamma(\mathcal{X},\mathcal{Y})\right\rVert_{\mathbb{K}}^{2}$ $=\underbrace{\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|{\mathcal{X}_{0}+\int_{0}^{t}\left(A\,\mathcal{X}_{s}+B\,\mathcal{Y}_{s}+b_{s}\right)\,\mathrm{d}s}\right|^{2}\right]}_{=:{\mathfrak{A}}}$ (3.60)
$\quad+\underbrace{\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\Bigg{|}{\mathbb{E}\left[G\mathcal{X}_{T}-\int_{t}^{T}\left(\hat{A}\,\mathcal{X}_{s}+\hat{B}\,\mathcal{Y}_{s}+\hat{b}_{s}\right)\,\mathrm{d}s\,\Big{|}\,{\mathcal{F}}_{t}\right]}\Bigg{|}^{2}\mathrm{d}t\right]}_{=:{\mathfrak{B}}}\,.$ (3.61)
For the first term, using the triangle inequality and Cauchy-Schwarz inequality we find the upper bound
${\mathfrak{A}}$ $\leq 2\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}|\mathcal{X}_{t}|^{2}\right]+2\,T\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\int_{0}^{t}\left|A\,\mathcal{X}_{s}+B\,\mathcal{Y}_{s}+b_{s}\right|^{2}\,\mathrm{d}s\right]$ (3.62)
$\leq 2\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}|\mathcal{X}_{t}|^{2}\right]+2\,T\,\mathbb{E}\left[\int_{0}^{T}\left|A\,\mathcal{X}_{s}+B\,\mathcal{Y}_{s}+b_{s}\right|^{2}\,\mathrm{d}s\right]$ (3.63)
$\leq 2\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}|\mathcal{X}_{t}|^{2}\right]+6\,T^{2}\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}|A|^{2}\,|\mathcal{X}_{t}|^{2}+|B|^{2}\,|\mathcal{Y}_{t}|^{2}+|b_{t}|^{2}\right]\,,$ (3.64)
which is finite because $(\mathcal{X},\mathcal{Y})\in\mathbb{K}$. For the second term, using Cauchy-Schwarz inequality we find the following upper bound
${\mathfrak{B}}$ $\leq 3\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathbb{E}\left[G\,\mathcal{X}_{T}\,\Big{|}\,{\mathcal{F}}_{t}\right]\right|^{2}\right]+3\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathbb{E}\left[\int_{0}^{T}\left(\hat{A}\,\mathcal{X}_{s}+\hat{B}\,\mathcal{Y}_{s}+\hat{b}_{s}\right)\,\mathrm{d}s\,\Big{|}\,{\mathcal{F}}_{t}\right]\right|^{2}\right]$ (3.65)
$\qquad+3\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\int_{0}^{t}\left(\hat{A}\,\mathcal{X}_{s}+\hat{B}\,\mathcal{Y}_{s}+\hat{b}_{s}\right)\,\mathrm{d}s\right|^{2}\right]\,.$ (3.66)
The third term above can be bound similarly to the bound for ${\mathfrak{A}}$. For the first two terms above, we use Doob’s martingale inequality to find that
$\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathbb{E}\left[G\,\mathcal{X}_{T}\,\Big{|}\,{\mathcal{F}}_{t}\right]\right|^{2}\right]+\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathbb{E}\left[\int_{0}^{T}\left(\hat{A}\,\mathcal{X}_{s}+\hat{B}\,\mathcal{Y}_{s}+\hat{b}_{s}\right)\,\mathrm{d}s\,\Big{|}\,{\mathcal{F}}_{t}\right]\right|^{2}\right]$ (3.67)
$\qquad\leq 4\,\mathbb{E}\left[\left|G\,\mathcal{X}_{T}\right|^{2}\right]+4\,\mathbb{E}\left[\left|\int_{0}^{T}\left(\hat{A}\,\mathcal{X}_{s}+\hat{B}\,\mathcal{Y}_{s}+\hat{b}_{s}\right)\,\mathrm{d}s\right|^{2}\right]$ (3.68)
$\qquad\leq 4\,|G|^{2}\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}|\mathcal{X}_{t}|^{2}\right]+4\,\mathbb{E}\left[\left|\int_{0}^{T}\left(\hat{A}\,\mathcal{X}_{s}+\hat{B}\,\mathcal{Y}_{s}+\hat{b}_{s}\right)\,\mathrm{d}s\right|^{2}\right]\,,$ (3.69)
from which we conclude that ${\mathfrak{B}}$ is finite. Thus, $\left\lVert\Gamma(\mathcal{X},\mathcal{Y})\right\rVert_{\mathbb{K}}^{2}<+\infty$ and it follows that $\Gamma(\mathcal{X},\mathcal{Y})\in\mathbb{K}$ because $\Gamma(\mathcal{X},\mathcal{Y})$ is adapted by construction.
Next, we show that for $T$ small enough $\Gamma$ is a contraction mapping. Let $(\mathcal{X},\mathcal{Y})\in\mathbb{K}$ and $(\tilde{\mathcal{X}},\tilde{\mathcal{Y}})\in\mathbb{K}$, it follows that
$\left\lVert\Gamma(\mathcal{X},\mathcal{Y})-\Gamma(\tilde{\mathcal{X}},\tilde{\mathcal{Y}})\right\rVert^{2}_{\mathbb{K}}$ (3.70)
$\quad=\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|{\int_{0}^{t}A\,\left(\mathcal{X}_{s}-\tilde{\mathcal{X}}_{s}\right)+B\,\left(\mathcal{Y}_{s}-\tilde{\mathcal{Y}}_{s}\right)\mathrm{d}s}\right|^{2}\right]$ (3.71)
$\quad\qquad+\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathbb{E}\left[G\,\left(\mathcal{X}_{T}-\tilde{\mathcal{X}}_{T}\right)-{\int_{t}^{T}\hat{A}\,\left(\mathcal{X}_{s}-\tilde{\mathcal{X}}_{s}\right)+\hat{B}\,\left(\mathcal{Y}_{s}-\tilde{\mathcal{Y}}_{s}\right)\mathrm{d}s}\,\Big{|}\,{\mathcal{F}}_{t}\right]\right|^{2}\right]\,.$ (3.72)
Using the triangle inequality, the Cauchy-Schwartz inequality, and Doob’s martingale inequality, we find that
$\left\lVert\Gamma(\mathcal{X},\mathcal{Y})-\Gamma(\tilde{\mathcal{X}},\tilde{\mathcal{Y}})\right\rVert^{2}_{\mathbb{K}}$ (3.73a)
$\qquad\leq 2\,T^{2}\,|A|^{2}\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathcal{X}_{t}-\tilde{\mathcal{X}}_{t}\right|^{2}\right]+2\,T^{2}\,|B|^{2}\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathcal{Y}_{t}-\tilde{\mathcal{Y}}_{t}\right|^{2}\right]$ (3.73b)
$\qquad\qquad+3\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathbb{E}\left[G\,\left(\mathcal{X}_{T}-\tilde{\mathcal{X}}_{T}\right)\,\Big{|}\,{\mathcal{F}}_{t}\right]\right|^{2}\right]$ (3.73c)
$\qquad\qquad+3\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathbb{E}\left[{\int_{0}^{T}\hat{A}\,\left(\mathcal{X}_{s}-\tilde{\mathcal{X}}_{s}\right)+\hat{B}\,\left(\mathcal{Y}_{s}-\tilde{\mathcal{Y}}_{s}\right)\mathrm{d}s}\,\Big{|}\,{\mathcal{F}}_{t}\right]\right|^{2}\right]$ (3.73d)
$\qquad\qquad+3\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\int_{0}^{t}\hat{A}\,\left(\mathcal{X}_{s}-\tilde{\mathcal{X}}_{s}\right)+\hat{B}\,\left(\mathcal{Y}_{s}-\tilde{\mathcal{Y}}_{s}\right)\mathrm{d}s\right|^{2}\right]$ (3.73e)
$\qquad\leq 2\,T^{2}\,|A|^{2}\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathcal{X}_{t}-\tilde{\mathcal{X}}_{t}\right|^{2}\right]+2\,T^{2}\,|B|^{2}\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathcal{Y}_{t}-\tilde{\mathcal{Y}}_{t}\right|^{2}\right]$ (3.73f)
$\qquad\qquad+12\,|G|^{2}\,\mathbb{E}\left[\left|\mathcal{X}_{T}-\tilde{\mathcal{X}}_{T}\right|^{2}\right]$ (3.73g)
$\qquad\qquad+12\,\mathbb{E}\left[\left|\int_{0}^{T}\hat{A}\,\left(\mathcal{X}_{s}-\tilde{\mathcal{X}}_{s}\right)+\hat{B}\,\left(\mathcal{Y}_{s}-\tilde{\mathcal{Y}}_{s}\right)\mathrm{d}s\right|^{2}\right]$ (3.73h)
$\qquad\qquad+6\,T^{2}\,|\hat{A}|^{2}\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathcal{X}_{t}-\tilde{\mathcal{X}}_{t}\right|^{2}\right]+6\,T^{2}\,|\hat{B}|^{2}\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathcal{Y}_{t}-\tilde{\mathcal{Y}}_{t}\right|^{2}\right]\,.$ (3.73i)
$\qquad\leq 2\,T^{2}\,|A|^{2}\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathcal{X}_{t}-\tilde{\mathcal{X}}_{t}\right|^{2}\right]+2\,T^{2}\,|B|^{2}\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathcal{Y}_{t}-\tilde{\mathcal{Y}}_{t}\right|^{2}\right]$ (3.73j)
$\qquad\qquad+12\,|G|^{2}\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathcal{X}_{t}-\tilde{\mathcal{X}}_{t}\right|^{2}\right]$ (3.73k)
$\qquad\qquad+24\,T^{2}\,|\hat{A}|^{2}\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathcal{X}_{t}-\tilde{\mathcal{X}}_{t}\right|^{2}\right]+24\,T^{2}\,|\hat{B}|^{2}\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathcal{Y}_{t}-\tilde{\mathcal{Y}}_{t}\right|^{2}\right]$ (3.73l)
$\qquad\qquad+6\,T^{2}\,|\hat{A}|^{2}\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathcal{X}_{t}-\tilde{\mathcal{X}}_{t}\right|^{2}\right]+6\,T^{2}\,|\hat{B}|^{2}\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathcal{Y}_{t}-\tilde{\mathcal{Y}}_{t}\right|^{2}\right]$ (3.73m)
$\qquad=T^{2}\,(12\,|G|^{2}+(2\,|A|^{2}+30\,|\hat{A}|^{2}))\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathcal{X}_{t}-\tilde{\mathcal{X}}_{t}\right|^{2}\right]$ (3.73n)
$\qquad\quad+T^{2}\,(2|B|^{2}+30|\hat{B}|^{2})\,\mathbb{E}\left[\sup_{t\in{\mathfrak{T}}}\left|\mathcal{Y}_{t}-\tilde{\mathcal{Y}}_{t}\right|^{2}\right]$ (3.73o)
Given the conditions on $|G|$ and $T$ it follows that $\Gamma$ is a contraction mapping and, by the Banach fixed point theorem in the Banach space $(\mathbb{K},||\cdot||_{\mathbb{K}})$, we have that there exists a unique fixed point. It is then easy to see that the fixed point is the solution to the FBSDE system in (3.51). ∎
The norms $|{A}|,|{B}|,|{\hat{A}}|,|{\hat{B}}|$ are easy to compute for specific examples;444 The expressions for $|{A}|,|{B}|$, and $|{\hat{B}}|$ are easy to obtain in general, but $|{\hat{A}}|$ is more involved. For example, for $|{A}|,|{B}|$, and $|{\hat{B}}|$, the general expressions are $|A|=\mathfrak{p}$, $|B|=\sqrt{3+\mathfrak{h}^{2}+\sqrt{5-2\,\mathfrak{h}^{2}+\mathfrak{h}^{4}}}/\sqrt{2}$, and $|\hat{B}|$ $=\max\Bigg{\{}\frac{\mathfrak{h}}{2\,b},\sqrt{4\,a^{2}\,\mathfrak{p}^{2}+\mathfrak{h}^{2}(1+\mathfrak{p}^{4})-\sqrt{-16\,a^{2}\,\mathfrak{h}^{2}\,\mathfrak{p}^{2}+(4\,a^{2}\,\mathfrak{p}^{2}+\mathfrak{h}^{2}\,(1+\mathfrak{p}^{4}))^{2}}}/(2\sqrt{2}\,a),$ (3.74) $\qquad\qquad\sqrt{4\,a^{2}\,\mathfrak{p}^{2}+\mathfrak{h}^{2}\,(1+\mathfrak{p}^{4})+\sqrt{-16\,a^{2}\,\mathfrak{h}^{2}\,\mathfrak{p}^{2}+(4\,a^{2}\,\mathfrak{p}^{2}+\mathfrak{h}^{2}\,(1+\mathfrak{p}^{4}))^{2}}}/(2\,\sqrt{2}\,a)\Bigg{\}}\,.$ (3.75) for example, in the numerical experiments below, $|A|=0$, $|B|=1.618$, $|\hat{A}|=1$, and $|\hat{B}|=0.5$.
4 Closed-form solution to the Nash equilibrium
Let $\ell:{\mathfrak{T}}\times\Omega\to\mathbb{R}^{3}$ and $\mathcal{P}:{\mathfrak{T}}\to\mathbb{R}^{3\times 3}$ be given by
${\ell}_{t}=\begin{pmatrix}g_{t}\\ h_{t}\\ f_{t}\end{pmatrix},\qquad\mathcal{P}_{t}=\begin{pmatrix}g^{B}_{t}&g^{I}_{t}&g^{Y}_{t}\\ h^{B}_{t}&h^{I}_{t}&h^{Y}_{t}\\ f^{B}_{t}&f^{I}_{t}&f^{Y}_{t}\end{pmatrix},$ (4.1)
and consider the ansatz
$\mathcal{Y}_{t}=\ell_{t}+\mathcal{P}_{t}\,\mathcal{X}_{t}\,.$ (4.2)
It follows that
$\mathrm{d}\mathcal{Y}_{t}$ $=\mathrm{d}\ell_{t}+\mathcal{P}^{\prime}_{t}\,\mathcal{X}_{t}\,\mathrm{d}t+\mathcal{P}_{t}\,\mathrm{d}\mathcal{X}_{t}$ (4.3)
$=\mathrm{d}\ell_{t}+\mathcal{P}^{\prime}_{t}\,\mathcal{X}_{t}\,\mathrm{d}t+\mathcal{P}_{t}\,\left(A\,\mathcal{X}_{t}+B\,\mathcal{Y}_{t}+b_{t}\right)\,\mathrm{d}t$ (4.4)
$=\mathrm{d}\ell_{t}+\mathcal{P}^{\prime}_{t}\,\mathcal{X}_{t}\,\mathrm{d}t+\mathcal{P}_{t}\,\left(A\,\mathcal{X}_{t}+B\,\left(\ell_{t}+\mathcal{P}_{t}\,\mathcal{X}_{t}\right)+b_{t}\right)\,\mathrm{d}t\,,$ (4.5)
which together with (3.51b) implies that
$\mathrm{d}\ell_{t}+\mathcal{P}^{\prime}_{t}\,\mathcal{X}_{t}\,\mathrm{d}t+\mathcal{P}_{t}\,\left(A\,\mathcal{X}_{t}+B\,\left(\ell_{t}+\mathcal{P}_{t}\,\mathcal{X}_{t}\right)+b_{t}\right)\,\mathrm{d}t=\left(\hat{A}\,\mathcal{X}_{t}+\hat{B}\,\left(\ell_{t}+\mathcal{P}_{t}\,\mathcal{X}_{t}\right)+\hat{b}_{t}\right)\,\mathrm{d}t+\sigma_{t}\,\mathrm{d}\mathcal{M}_{t}\,.$ (4.6)
Rearranging the above expression and collecting terms with $\mathcal{X}$ we obtain that
$\begin{split}0&=\left\{\mathrm{d}\ell_{t}+\left[\mathcal{P}_{t}\left(B\,\ell_{t}+b_{t}\right)-\hat{B}\,\ell_{t}-\hat{b}_{t}\right]\mathrm{d}t-\sigma_{t}\,\mathrm{d}\mathcal{M}_{t}\right\}\\ &\qquad+\left\{\mathcal{P}^{\prime}_{t}+\mathcal{P}_{t}\,A+\mathcal{P}_{t}\,B\,\mathcal{P}_{t}-\hat{A}-\hat{B}\,\mathcal{P}_{t}\right\}\mathcal{X}_{t}\,\mathrm{d}t\,.\end{split}$ (4.7)
Thus, showing existence and uniqueness of the solution to the FBSDE in (3.51) boils down to showing existence and uniqueness of the non-Hermitian matrix Riccati differential equation
00 0 $=\mathcal{P}^{\prime}_{t}+\mathcal{P}_{t}\,A+\mathcal{P}_{t}\,B\,\mathcal{P}_{t}-\hat{A}-\hat{B}\,\mathcal{P}_{t}\,,$ (4.8a)
$\mathcal{P}_{T}$ $=G\,,$ (4.8b)
and the linear BSDE
00 0 $=\mathrm{d}\ell_{t}+\left[\mathcal{P}_{t}\left(B\,\ell_{t}+b_{t}\right)-\hat{B}\,\ell_{t}-\hat{b}_{t}\right]\mathrm{d}t-\sigma_{t}\,\mathrm{d}\mathcal{M}_{t}\,,$ (4.9a)
$\ell_{T}$ $=0\,.$ (4.9b)
Subsection 4.1 shows existence and uniqueness of a solution to (4.8) when $\mathfrak{p}=r^{B}=0$ under no further restrictions of model parameters, and Subsection 4.2 shows existence and uniqueness when $\mathfrak{p}=0$ under four additional mild restrictions on model parameters. Lastly, as mentioned before, when $\mathfrak{p}>0$ existence and uniqueness of the FBSDE follow from the assumption that $T$ is small enough; see Theorem 3.8 for the bound on the time horizon $T$.
4.1 No resilience and no running penalty for the broker
Next, when the values of model parameters $\mathfrak{p}$ and $r^{B}$ are zero, we show existence and uniqueness of a solution to (4.8) with no additional constraints on the values of the remaining model parameters. Let
${C}=\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&1\end{pmatrix}\,,\qquad{D}=\begin{pmatrix}-1&0&0\\ 0&-1&0\\ 0&0&0\end{pmatrix}\,,$ (4.10)
and use Theorem 2.3 in Freiling et al., (2000) with the matrices ${C}$ and ${D}$ above. More precisely, using the notation of Freiling et al., (2000), we have that $B_{11}=A$, $B_{12}=B$, $B_{21}=\hat{A}$, and $B_{22}=\hat{B}$. Then, it follows that for our choice of $C$ and $D$,
${C}+{D}\,G+G^{\intercal}\,{D}^{\intercal}=\begin{pmatrix}\frac{2\,\varphi}{a}&0&0\\ 0&\frac{2\,\psi}{b}&0\\ 0&0&1\end{pmatrix}>0\,,$ (4.11)
which is positive definite because $\varphi,\psi>0$. Next, matrix ${L}$ in Theorem 2.3 of Freiling et al., (2000) is given by
${L}=\begin{pmatrix}{C}\,B_{11}+{D}\,B_{21}&{C}\,B_{12}+B_{11}^{\intercal}{D}+{D}\,B_{22}\\ 0&B_{12}^{\intercal}{D}\end{pmatrix}\,,$ (4.12)
and a short calculation (using Sylvester’s criterion) shows that ${L}+{L}^{\intercal}\leq 0$. Similar to part II of the proof of Theorem 3.5 in Casgrain and Jaimungal, (2020), the solution is bounded because the solution exists and is continuous in $[0,T]$; therefore, from Theorem 3.1 in Freiling, (2002), the unique solution takes the form
$\mathcal{P}_{t}=\mathcal{T}_{t}\,{R}^{-1}_{t}\,,$ (4.13)
where the pair ${R}_{t},\mathcal{T}_{t}$ is the unique solution to the linear system of differential equations
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{pmatrix}{R}_{t}\\ \mathcal{T}_{t}\end{pmatrix}=\begin{pmatrix}B_{11}&B_{12}\\ B_{21}&B_{22}\end{pmatrix}\,\begin{pmatrix}{R}_{t}\\ \mathcal{T}_{t}\end{pmatrix}\,,\qquad\begin{pmatrix}{R}_{T}\\ \mathcal{T}_{T}\end{pmatrix}=\begin{pmatrix}I\\ G\end{pmatrix}\,.$ (4.14)
4.2 No resilience
[…]
4.3 The solution to the BSDE
Next, we use the solution to the non-Hermitian matrix Riccati differential equation to find a solution to the BSDE (4.9) that $\ell$ must satisfy. Write (4.9) as
$\mathrm{d}{\ell}_{t}=\left({a}_{t}+\Lambda_{t}\,{\ell}_{t}\right)\mathrm{d}t+{\sigma}_{t}\,\mathrm{d}\mathcal{M}_{t}\,,$ (4.15)
where $a_{t}=\hat{b}_{t}-\mathcal{P}_{t}\,b_{t}$ and $\Lambda_{t}=\hat{B}-\mathcal{P}_{t}\,B$, and let ${\zeta}_{t}$ be the unique solution to
$\mathrm{d}{\zeta}_{t}=-{\zeta}_{t}\,{\Lambda}_{t}\,\mathrm{d}t\,,$ (4.16)
with initial condition ${\zeta}_{0}=I\in\mathbb{R}^{2\times 2}$. Then,
$\mathrm{d}\left({\zeta}_{t}\,{\ell}_{t}\right)$ $={\ell}_{t}\,\mathrm{d}{\zeta}_{t}+{\zeta}_{t}\,\mathrm{d}{\ell}_{t}$ (4.17)
$=-{\ell}_{t}\,{\zeta}_{t}\,{\Lambda}_{t}\,\mathrm{d}t+{\zeta}_{t}\,\left({a}_{t}+\Lambda_{t}\,{\ell}_{t}\right)\mathrm{d}t-{\zeta}_{t}\,{\sigma}_{t}\,\mathrm{d}\mathcal{M}_{t}$ (4.18)
$={\zeta}_{t}\,{a}_{t}\,\mathrm{d}t-{\zeta}_{t}\,{\sigma}_{t}\,\mathrm{d}\mathcal{M}_{t}\,,$ (4.19)
which implies that
$-{\zeta}_{t}\,{\ell}_{t}$ $=\int_{t}^{T}{\zeta}_{u}\,{a}_{u}\,\mathrm{d}u+\int_{t}^{T}{\zeta}_{u}\,{\sigma}_{u}\,\mathrm{d}\mathcal{M}_{u}\,,$ (4.20)
because ${\zeta}_{T}\,{\ell}_{T}=0$. Thus, by taking conditional expectations, the solution to ${\ell}_{t}$ is
${\ell}_{t}$ $=-\mathbb{E}\left[\int_{t}^{T}{\zeta}^{-1}_{t}\,{\zeta}_{u}\,{a}_{u}\,\mathrm{d}u\,\big{|}\,{\mathcal{F}}_{t}\right]\,.$ (4.21)
We conclude that the Nash equilibrium $\nu^{*},\eta^{*}$, which are the first two components of $\mathcal{Y}$, is given by
$\mathcal{Y}_{t}={\ell}_{t}+\mathcal{P}_{t}\,\mathcal{X}_{t}\,,$ (4.22)
where $\mathcal{X}$ is the unique solution to the forward equation (3.51a).
5 Numerical results
In this section, we study the optimal trading strategies and the Nash equilibrium between the broker and the informed trader. We discretise the trading window $[0,T]$, with $T=1$, into 10,000 steps and perform one million simulations. The terminal penalty parameters are $\psi=\phi=1$, the instantaneous price impact parameters are $b=1\times 10^{-3}$ and $a=1.2\times 10^{-3}$, the running penalties are $r^{I}=r^{B}=0$, and the transient price impact parameters are $\mathfrak{h}=10^{-3}$, and $\mathfrak{p}=0$. The midprice starts at $S_{0}=100$ and $M^{S}_{t}=\sigma^{S}\,W^{S}_{t}$, where $\sigma^{S}=1$ and $W^{S}_{t}$ is a Brownian motion. The signal $\alpha_{t}$ follows the OU process $\mathrm{d}\alpha_{t}=-\kappa^{\alpha}\,\alpha_{t}\,\mathrm{d}t+\sigma^{\alpha}\,\mathrm{d}W^{\alpha}_{t}$, where $\kappa^{\alpha}=5$, $\sigma^{\alpha}=1$, $\alpha_{0}=0$, and $W^{\alpha}_{t}$ a Brownian motion. Finally, the trading speed of the uninformed trader follows the OU process $\mathrm{d}\xi_{t}=-\kappa^{\xi}\,\xi_{t}\,\mathrm{d}t+\sigma^{\xi}\,\mathrm{d}W^{\xi}_{t}$, where $\kappa^{\xi}=15$, $\sigma^{\xi}=100$, $\xi_{0}=0$, and $W^{\xi}$ is a Brownian motion. The three Brownian motions $W^{\xi}$, $W^{S}$, and $W^{\alpha}$ are independent, for simplicity. Next, we provide a closed-form representation of (4.21), given by
${\ell}_{t}=\begin{pmatrix}g_{t}\\ h_{t}\\ f_{t}\end{pmatrix}=\begin{pmatrix}g^{1}(t)\,\alpha_{t}+g^{2}(t)\,\xi_{t}\\ h^{1}(t)\,\alpha_{t}+h^{2}(t)\,\xi_{t}\\ f^{1}(t)\,\alpha_{t}+f^{2}(t)\,\xi_{t}\\ \end{pmatrix}\,,$ (5.1)
where $g^{1,2}$, $h^{1,2}$, and $f^{1,2}$ are deterministic functions characterised by the unique solution to the linear system of ODEs
00 0 $=\frac{\mathrm{d}g^{1}_{t}}{\mathrm{d}t}+g^{B}_{t}\,\left(g^{1}_{t}-h^{1}_{t}\right)+g^{I}_{t}\,h^{1}_{t}+g^{1}_{t}\,g^{Y}_{t}\,\mathfrak{h}-g^{1}_{t}\,\kappa^{\alpha}+\frac{1}{2\,a}+\frac{h^{1}_{t}\,\mathfrak{h}}{2\,a}\,,$ (5.2a)
00 0 $=\frac{\mathrm{d}g^{2}_{t}}{\mathrm{d}t}-g^{B}_{t}+g^{B}_{t}\,\left(g^{2}_{t}-h^{2}_{t}\right)+g^{I}_{t}\,h^{2}_{t}+g^{2}_{t}\,g^{Y}_{t}\,\mathfrak{h}-g^{2}_{t}\,\kappa^{\xi}+\frac{\mathfrak{h}}{2\,a}+\frac{h^{2}_{t}\,\mathfrak{h}}{2\,a}\,,$ (5.2b)
00 0 $=\frac{\mathrm{d}h^{1}_{t}}{\mathrm{d}t}+h^{B}_{t}\,\left(g^{1}_{t}-h^{1}_{t}\right)+h^{1}_{t}\,h^{I}_{t}+g^{1}_{t}\,h^{Y}_{t}\,\mathfrak{h}-h^{1}_{t}\,\kappa^{\alpha}+\frac{1}{2\,b}+\frac{g^{1}_{t}\,\mathfrak{h}}{2\,b}\,,$ (5.2c)
00 0 $=\frac{\mathrm{d}h^{2}_{t}}{\mathrm{d}t}-h^{B}_{t}+h^{B}_{t}\left(g^{2}_{t}-h^{2}_{t}\right)+h^{2}_{t}\,h^{I}_{t}+g^{2}_{t}\,h^{Y}_{t}\,\mathfrak{h}-h^{2}_{t}\,\kappa^{\xi}+\frac{g^{2}_{t}\,\mathfrak{h}}{2\,b}\,,$ (5.2d)
00 0 $=\frac{\mathrm{d}f^{1}_{t}}{\mathrm{d}t}+f^{B}_{t}\left(g^{1}_{t}-h^{1}_{t}\right)+f^{I}_{t}\,h^{1}_{t}+f^{Y}_{t}g^{1}_{t}\,\mathfrak{h}-f^{1}_{t}\,\kappa^{\alpha}\,,$ (5.2e)
00 0 $=\frac{\mathrm{d}f^{2}_{t}}{\mathrm{d}t}-f^{B}_{t}+f^{B}_{t}\left(g^{2}_{t}-h^{2}_{t}\right)+f^{I}_{t}\,h^{2}_{t}+f^{Y}_{t}\,g^{2}_{t}\,\mathfrak{h}-f^{2}_{t}\,\kappa^{\xi}\,,$ (5.2f)
with terminal conditions $g^{1,2}_{T}=0$, $h^{1,2}_{T}=0$, and $f^{1,2}_{T}=0$. Given that $\mathfrak{p}=0$, we require only $g^{B},g^{I},h^{B},h^{I}:[0,T]\to\mathbb{R}$ as then, $f^{B},f^{I},f^{Y},g^{Y},h^{Y}$ do not enter the optimal strategies of the broker and informed traders.
Figure 1 shows a sample path for the trading signal $\alpha$, the optimal trading speeds of the informed trader ($\eta^{*}$) and the broker ($\nu^{*}$) together with their inventories $Q^{I,\eta^{*}}$ and $Q^{B,\nu^{*}}$. Additionally, we include the $5\%$ and $95\%$ quantiles through time for these quantities.
Figure 1: Selected sample paths for the main processes that play a role in the Nash equilibrium.
Next, we provide a comparison between the resulting inventory paths as the transient impact rate of decay parameter $\mathfrak{p}$ varies. To this end, Figure 2 shows a single sample path of the broker’s and informed trader’s inventory together with the $5\%$ and $95\%$ quantile bands through time for four choices of the decay coefficient $\mathfrak{p}$ with $\mathfrak{h}=10$ and all parameters the same as in the previous example. As seen in the figure, as $\mathfrak{p}$ increases, the broker’s inventory paths appear to be more concentrated around zero, while the informed trader’s inventory initially becomes more concentrated with increasing $\mathfrak{p}$, it appears to widen for the largest value of $\mathfrak{p}=16$.
Figure 2: Sample path together with 5% and 95% quantile bands for the trajectory of the inventory of the broker (left panel) and the informed trader (right panel). We show the plots associated the values $\mathfrak{p}\in\{0,4,8,16\}$ of the decay coefficient.
Finally, we compare the two-stage optimisation in Cartea and Sánchez-Betancourt, (2022) with the Nash equilibrium we find. We employ the same model parameters as above and set the ambiguity aversion parameters in their paper to $0.001$. We obtain 100,000 sample paths for the mark-to-market of the broker and informed trader when following the strategies from the Nash equilibrium and those from the two-stage optimisation, all using the same set of Brownian paths. Figure 3 shows the resulting kernel density estimator for all six pairwise quantities, where Nash-informed (Nash-broker) refers to the mark-to-market of the informed (broker) trader when following the Nash-equilibrium strategy, and $\Delta$ informed ($\Delta$ broker) refers to the the difference between the mark-to-market when following the Nash equilibrium and the mark-to-market when following the two-stage optimisation strategy for the informed trader (broker).
Figure 3: Contour plots of the kernel density estimator for the Nash equilibrium found here versus the two-stage optimisation in Cartea and Sánchez-Betancourt, (2022). We report the correlation $\rho$ between the two variables.
The mean mark-to-market (MtM) of the informed trader increases by roughly 4.5% when going from the two-stage optimisation in Cartea and Sánchez-Betancourt, (2022) to the Nash equilibrium, whereas the broker decreases their mean MtM by 5.5%. A $t$-test shows, with 99% confidence, that the average MtM of the samples generated by both the Nash and two-stage optimisation are statistically different (for both the informed trader and the broker). This difference provides an upper bound for how profitability moves between players when information becomes publicly available. Thus, in our model (for the parameters above), the maximum improvement that the informed trader can obtain is roughly 5% which is largely paid by the broker; note that this is four orders of magnitude higher than transaction costs in foreign exchange markets. [^5] Transaction costs in spot foreign exchange are roughly \$3/\$ M; see Cartea and Sánchez-Betancourt, (2023). Our results provide insights into what happens in broker-client relationships in markets where the broker does not benefit from anonymity. Such perfect information trading environments are a reality in the decentralised finance space – in particular in automated market maker trading venues.
6 Acknowledgements
SJ would like to thank the National Sciences and Engineering Research council of Canada for partially supporting this research through grants RGPIN-2018-05705 and RGPIN-2024-04317.